### v1

# Fully discrete error analysis of finite element discretizations of time-dependent Stokes equations in a stream-function formulation

Abstract
In this paper we establish best approximation type estimates for the fully discrete Galerkin solutions of time-dependent Stokes problem using the stream-function formulation. For the time discretization we use the discontinuous Galerkin method of arbitrary degree, whereas we present the space discretization in a general framework. This makes our result applicable for a wide variety of space discretization methods, provided some Galerkin orthogonality conditions are satisfied. As an example, conformal $C^{1}$ and $C^{0}$ interior penalty methods are covered by our analysis. The results do not require any additional regularity assumptions beyond the natural regularity given by the domain and data and can be used for optimal control problems.
keywords:
Stokes problem, time-dependent Stokes, parabolic problems, steamline formulation, finite elements, discontinuous Galerkin, error estimates, best approximation, fully discrete
AMS:
(MSC2020) 76M10, 35G16, 65M15, 65M60, 65N30
1 Introduction
In this paper we consider the following time-dependent Stokes problem with no-slip boundary conditions,
$\partial_{t}\mathbf{u}-\Delta\mathbf{u}+\nabla p$ $=\mathbf{g}\quad$ $\text{in }I\times\Omega,$ (1)
$\nabla\cdot\mathbf{u}$ $=0\quad$ $\text{in }I\times\Omega,$
$\mathbf{u}$ $=\mathbf{0}\quad$ $\text{on }I\times\partial\Omega,$
$\mathbf{u}(0)$ $=\mathbf{u}_{0}$ $\text{in }\Omega.$
We assume that $\Omega\subset\mathbb{R}^{2}$, is a bounded, polygonal and simply connected domain, $T>0$ and $I=(0,T]$. The finite element approximation of Stokes and Navier-Stokes problems is a mature and well researched topic due to its importance for understanding and computing fluid problems. The literature is immense, we only would like to mention the monographs [12, 15, 31].
The numerical, and in particular finite element approximation of the problem, has a long history and the research is still very active. There are many challenges one faces in order to construct proper finite element spaces. It is well-known that the spaces for the discrete velocity $\mathbf{u}_{h}$ and the discrete pressure $p_{h}$ can not be taken arbitrary and the corresponding finite element spaces must satisfy the inf-sup condition. In addition, the continuity equation, i.e., $\nabla\cdot\mathbf{u}=0$, is not trivial to satisfy on the discrete level pointwise, and usually requires either high order methods [4, 14, 29] or very special constructions [19, 20]. The time discretization poses new challenges. For example, on non-matching meshes, a discretely divergence free function on the previous mesh is not discretely divergence free on the next mesh and needs to be projected, which leads to oscillations in the pressure. All these challenges are well-known and are addressed in more or less satisfactory fashion and constitute the standard part of the finite element theory. Recently, the finite element theory and practices were looked in a new light and many standard finite element approximation schemes faced some criticism. The main objection of many standard finite element methods, like Taylor-Hood or MINI elements, are the facts that the discrete pressure and velocity are coupled and the error estimates for the velocity contain a pressure term. As a result, in some situations, the perturbations of the source term by a gradient field is not completely absorbed by the discrete pressure alone and may lead to a large error for the computed velocity [21]. This observation had led recently to a development of various pressure robust methods [2, 3, 26].
An old two dimensional approach, that checks all the above mentioned points is the streamline function method where the velocity $\mathbf{u}$ is taken as vectorial curl of a scalar function $\psi$
$\mathbf{u}=\operatorname{\mathbf{curl}}(\psi):=(\partial_{2}\psi,-\partial_{1}\psi)^{T}.$ (2)
Its dual operator is given by the scalar curl of a vector field $\mathbf{v}$
$\operatorname{curl}(\mathbf{v})=\partial_{2}\mathbf{v}_{1}-\partial_{1}\mathbf{v}_{2}.$
For $\mathbf{v}\in H^{1}_{0}(\Omega)^{2},\psi\in H^{1}_{0}(\Omega)$, there hold the following identities, to be understood in $H^{-1}(\Omega)^{2}$ and $H^{-1}(\Omega)$, respectively:
$\operatorname{\mathbf{curl}}(\operatorname{curl}(\mathbf{v}))$ $=(\partial_{2}^{2}\mathbf{v}_{1}-\partial_{2}\partial_{1}\mathbf{v}_{2},-\partial_{1}\partial_{2}\mathbf{v}_{1}+\partial_{1}^{2}\mathbf{v}_{2})^{T}=\Delta\mathbf{v}-\nabla(\text{div }\mathbf{v})$ (3)
$\operatorname{curl}(\operatorname{\mathbf{curl}}(\psi))$ $=\Delta\psi.$ (4)
For velocities $\mathbf{u}$ satisfying the form (2), it clearly holds $\text{div }\mathbf{u}=0$, and taking formally the negative scalar curl of the equation (1), we find that the scalar function $\psi$ satisfies the following fourth order equation with $f=-\operatorname{curl}(\mathbf{g})$
$-\partial_{t}\Delta\psi+\Delta^{2}\psi$ $=f\quad$ $\text{in }I\times\Omega,$ (5)
$\psi=\partial_{n}\psi$ $=0\quad$ $\text{on }I\times\partial\Omega,$
$\psi(0)$ $=\psi_{0}$ $\text{in }\Omega,$
where $\psi_{0}$ is the solution to
$-\Delta\psi_{0}$ $=-\operatorname{curl}(\mathbf{u}_{0})\quad$ $\text{in }\Omega,$ (6)
$\psi_{0}$ $=0\quad$ $\text{on }\partial\Omega.$
For simply connected $\Omega$, the equations (1) and (5) are equivalent, for a proof, we refer to [17]. Although the stationary problem is well investigated, (cf. Parts 4 and 5 in [18]), it seems there is no clean theory of fully discrete finite element error estimates for Galerkin methods for the time dependent problem. For example, a relatively recent paper [1] treated backward Euler for virtual elements and derived some fully discrete error estimates in $L^{\infty}(H^{1})$ norm. However, besides treating only lowest order discretization the error analysis presented there is rather messy and involves unnecessary terms, like time derivatives for the estimation of the spacial error, as well as artificial assumptions on $\partial_{tt}\psi$ and $\mathbf{g}\in L^{\infty}(I;H^{r-1}(\Omega)^{2})$ for some $r>\frac{1}{2}$.
In this paper, we take a first step in the direction of developing an elegant theory for the fully discrete error estimates similarly to the one for the second order parabolic problems [23, 24]. We establish a best approximation type error estimate for discontinuous Galerkin discretizations in time of arbitrary degree and finite element discretization in space (that include conformal $C^{1}$ elements as well as $C^{0}$ interior penalty methods), namely
$\lVert{\rm NewA}(\psi-\psi_{{k}h})\rVert_{L^{2}(I\times\Omega)}\leq C\Big{(}\lVert{\rm NewA}(\psi-{\chi})\rVert_{L^{2}(I\times\Omega)}+\lVert{\rm NewA}(\psi-R_{h}\psi)\rVert_{L^{2}(I\times\Omega)}+\lVert{\rm NewA}(\psi-\pi_{k}\psi)\rVert_{L^{2}(I\times\Omega)}\Big{)},$
where $\psi_{kh}$ is the fully discrete Galerkin solution for (5), $R_{h}$ is the Galerkin projection of the stationary biharmonic problem and $\pi_{k}$ is a certain time projection, and $\chi$ is an arbitrary element from the finite dimensional space. Under low regularity assumptions, i.e. $f\in L^{2}(I;H^{-1}(\Omega))$ and $\psi_{0}\in H^{2}_{0}(\Omega)$, using the approximation properties of the above mentioned projections we can establish the following optimal convergence result
$\lVert{\rm NewA}(\psi-\psi_{kh})\rVert_{L^{2}(I\times\Omega)}\leq C(k+h^{2})\Big{(}\|f\|_{L^{2}(I;H^{-1}(\Omega))}+\|\psi_{0}\|_{H^{2}_{0}(\Omega)}\Big{)}$ (7)
on convex polygonal domains. The minimal assumptions we put on the data are satisfied by the usual assumptions on data for the time-dependent Stokes problem, namely $\mathbf{g}\in L^{2}(I;\mathbf{H})$ and $\mathbf{u}_{0}\in H^{1}_{0}(\Omega)^{2}$ with $\nabla\cdot\mathbf{u}_{0}=0$. For a precise introduction of the space $\mathbf{H}$, see Subsection 2.3. Indeed we will argue later, see (11) and Lemma 4, that under those assumptions it holds $f\in L^{2}(I;H^{-1}(\Omega))$ and $\psi_{0}\in H^{2}_{0}(\Omega)$. As a result, taking $\mathbf{u}_{kh}=\operatorname{\mathbf{curl}}(\psi_{kh})$ we obtain for the velocity solution of (1)
$\lVert\mathbf{u}-\mathbf{u}_{kh}\rVert_{L^{2}(I\times\Omega)}$ $=\lVert\operatorname{\mathbf{curl}}(\psi-\psi_{kh})\rVert_{L^{2}(I\times\Omega)}=\lVert{\rm NewA}(\psi-\psi_{kh})\rVert_{L^{2}(I\times\Omega)}$ (8)
$\leq C(k+h^{2})\left(\lVert\mathbf{g}\rVert_{L^{2}(I\times\Omega)}+\lVert\mathbf{u}_{0}\rVert_{H^{1}_{0}(\Omega)}\right).$
As we have formulated our main result in a best approximation type form, higher order estimates are also possible, assuming higher regularity of $\psi$. The remainder of the paper is structured as follows: in Section 2, we summarize some results of the continuous Stokes problem, and establish some results regarding the associated stationary and time-dependent biharmonic equations. In Section 3 we introduce the space and time discretizations, keeping the discussion of the space discretization very short and general, and establish the first stability result for the fully discrete problem. We proceed to prove the best approximation type error estimate in Section 4, with our main result being Theorem 23. After this, we discuss the specific orders of convergence for special choices of the space discretization. We conclude with some numerical experiments in Section 5.
2 Continuous problem
2.1 Notation
We will use the standard notation for Lebesgue and Hilbert spaces. Most of our analysis will be $L^{2}$ based and we will denote the space and space-times inner-products by
$\left(f,g\right)_{\Omega}=\int_{\Omega}fg\,dx\quad\text{and}\quad(f,g)_{I\times\Omega}=\int_{I}(f,g)_{\Omega}dt,$
respectively, with the corresponding norms
$\|f\|^{2}_{\Omega}=\int_{\Omega}|f|^{2}\,dx\quad\text{and}\quad\|f\|^{2}_{I\times\Omega}=\int_{I}\|f\|^{2}_{\Omega}dt.$
When the situation is less clear we will use the standard notation for the norms, $L^{2}(\Omega)$, $H^{2}(\Omega)$ and etc. We will denote by $\langle\cdot,\cdot\rangle$ a dual pairing between a function space and its dual.
2.2 Function spaces
We define the following functional spaces.
$H^{1}_{0}(\Omega)$ $=\{v\in H^{1}(\Omega)\ :\ v=0\quad\text{on $\partial\Omega$}\},$
$H^{2}_{0}(\Omega)$ $=\{v\in H^{2}(\Omega)\ :\ v=\partial_{n}v=0\quad\text{on $\partial\Omega$}\},$
$H^{-2}(\Omega)$ $=\left(H^{2}_{0}(\Omega)\right)^{*},\quad H^{-1}(\Omega)=\left(H^{1}_{0}(\Omega)\right)^{*}.$
Lemma 1.
On the function space $H^{2}_{0}(\Omega)$, the norm $\lVert v\rVert_{H^{2}_{0}(\Omega)}:=\lVert\Delta v\rVert_{L^{2}(\Omega)}$ is equivalent to $\lVert v\rVert_{H^{2}(\Omega)}$.
Proof.
The proof follows by the integration by parts. ∎
2.3 Results on the time-dependent Stokes problem
For a better comparison of the stream function formulation to the classic formulation of the Stokes problem, let us recall some well known results. Most commonly, the Stokes equations on the continuous level are discussed in divergence free spaces, to which end we introduce the notations
$\mathbf{V}:=\{\mathbf{v}\in H^{1}_{0}(\Omega)^{2}:\nabla\cdot\mathbf{v}=0\}\quad\text{ and }\quad\mathbf{H}:=\{\mathbf{v}\in L^{2}(\Omega)^{2}:\nabla\cdot\mathbf{v}=0\land\mathbf{n}\cdot\mathbf{v}=0\},$
where by $\mathbf{n}\cdot\mathbf{v}$ we denote the normal trace of $\mathbf{v}$. It is straightforward to verify the following result
Lemma 2.
On the space $H^{1}_{0}(\Omega)^{2}$, the following norms are equivalent:
$\lVert\mathbf{v}\rVert_{H^{1}_{0}(\Omega)}\cong\lVert{\rm NewA}\cdot\mathbf{v}\rVert_{L^{2}(\Omega)}+\lVert\operatorname{curl}(\mathbf{v})\rVert_{L^{2}(\Omega)}.$
Especially, $\lVert\operatorname{curl}(\mathbf{v})\rVert_{L^{2}(\Omega)}$ induces a norm on $\mathbf{V}$ equivalent to the $H^{1}_{0}(\Omega)$ norm.
Proof.
For $\mathbf{v}\in C^{\infty}_{0}(\Omega)^{2}$ we obtain using the definition of $\nabla\cdot$ and $\operatorname{curl}$, as well as integration by parts
$\|\nabla\cdot\mathbf{v}\|_{L^{2}(\Omega)}^{2}+\|\operatorname{curl}(\mathbf{v})\|_{L^{2}(\Omega)}^{2}$ $=\int_{\Omega}(\partial_{1}\mathbf{v}_{1}+\partial_{2}\mathbf{v}_{2})^{2}+(\partial_{1}\mathbf{v}_{2}-\partial_{2}\mathbf{v}_{1})^{2}\,dx$
$=\int_{\Omega}\sum_{i,j=1}^{2}(\partial_{i}\mathbf{v}_{j})^{2}+2\partial_{1}\mathbf{v}_{1}\partial_{2}\mathbf{v}_{2}-2\partial_{1}\mathbf{v}_{2}\partial_{2}\mathbf{v}_{1}\,dx$
$=\int_{\Omega}\nabla\mathbf{v}:\nabla\mathbf{v}-2\mathbf{v}_{1}\partial_{1}\partial_{2}\mathbf{v}_{2}+2(\partial_{2}\partial_{1}\mathbf{v}_{2})\mathbf{v}_{1}\,dx$
$=\int_{\Omega}\nabla\mathbf{v}:\nabla\mathbf{v}\,dx=\|\mathbf{v}\|_{H^{1}_{0}(\Omega)}^{2}.$
The statements of the lemma then follow from density of $C^{\infty}_{0}(\Omega)$ in $H^{1}_{0}(\Omega)$, and the definition of $\mathbf{V}$. ∎
Using these spaces, the stationary Stokes operator $A\colon D(A)\to\mathbf{H}$ is given by
$\left(A\mathbf{u},\mathbf{v}\right)_{\Omega}=\left(\nabla\mathbf{u},\nabla\mathbf{v}\right)_{\Omega}\quad\text{ for all }\mathbf{v}\in\mathbf{V},$ (9)
where $D(A):=\{\mathbf{v}\in\mathbf{V}:\Delta\mathbf{v}\in L^{2}(\Omega)^{2}\}$. The operators $\operatorname{curl},\operatorname{\mathbf{curl}}$ lead to the following isomorphisms between scalar valued Sobolev spaces and divergence free vector spaces, see [16, Lemma 2]:
$\operatorname{\mathbf{curl}}\colon$ $H^{1}_{0}(\Omega)\to\mathbf{H},$ and $\operatorname{\mathbf{curl}}\colon$ $H^{2}_{0}(\Omega)\to\mathbf{V},$ (10)
$\operatorname{curl}\colon$ $\mathbf{H}\to H^{-1}(\Omega),$ and $\operatorname{curl}\colon$ $\mathbf{V}^{*}\to H^{-2}(\Omega).$ (11)
The time-dependent Stokes equations are described by the spaces
$\mathbb{V}$ $:=L^{2}(I;\mathbf{V})\cap H^{1}(I;\mathbf{V}^{*})\hookrightarrow C(\bar{I};\mathbf{H}),$
$\mathbb{W}$ $:=L^{2}(I;\mathbf{V}\cap H^{2}(\Omega)^{2})\cap H^{1}(I;\mathbf{H})\hookrightarrow C(\bar{I};\mathbf{V}).$
The weak formulation of the time-dependent Stokes equations reads: Find $\mathbf{u}\in\mathbb{V}$ such that $\mathbf{u}(0)=\mathbf{u}_{0}$ and
$\left\langle\partial_{t}\mathbf{u},\mathbf{v}\right\rangle_{I\times\Omega}+\left(\nabla\mathbf{u},\nabla\mathbf{v}\right)_{I\times\Omega}=\left\langle\mathbf{g},\mathbf{v}\right\rangle_{I\times\Omega}\quad\forall\mathbf{v}\in L^{2}(I;\mathbf{V}),$ (12)
where $\left\langle\cdot,\cdot\right\rangle_{I\times\Omega}$ denotes the duality pairing between $L^{2}(I;\mathbf{V})$ and $L^{2}(I;\mathbf{V}^{*})$. There holds the following result TO ADD:
Proposition 3.
Let $\mathbf{g}\in L^{2}(I;V^{*})$ and $\mathbf{u}_{0}\in\mathbf{H}$, then (12) has a unique weak solution $\mathbf{u}\in\mathbb{V}$ with
$\lVert\mathbf{u}\rVert_{\mathbb{V}}\leq C(\lVert\mathbf{g}\rVert_{L^{2}(I;\mathbf{V}^{*})}+\lVert\mathbf{u}_{0}\rVert_{\mathbf{H}}).$
If $\mathbf{g}\in L^{2}(I;L^{2}(\Omega)^{2})$ and $\mathbf{u}_{0}\in\mathbf{V}$, then $\mathbf{u}$ additionally satisfies
$\lVert\partial_{t}\mathbf{u}\rVert_{L^{2}(I;\mathbf{H})}+\lVert A\mathbf{u}\rVert_{L^{2}(I;\mathbf{H})}\leq C(\lVert\mathbf{g}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\mathbf{u}_{0}\rVert_{\mathbf{V}}),$
where $A$ is the Stokes operator defined in (9). If moreover $\Omega$ is convex, then $\mathbf{u}\in\mathbb{W}$ with a bound
$\lVert\mathbf{u}\rVert_{\mathbb{W}}\leq C(\lVert\mathbf{g}\rVert_{L^{2}(I;L^{2}(\Omega))}+\lVert\mathbf{u}_{0}\rVert_{\mathbf{V}}).$
We have already stated, that the initial data $\psi_{0}$ for the stream-function formulation and the data $\mathbf{u}_{0}$ for the standard formulation have to satisfy the relation (6). In the following lemma, we establish that the regularity assumptions posed on $\psi_{0}$ are actually natural and hold true, if the corresponding standard assumptions are posed on $\mathbf{u}_{0}$.
Lemma 4.
Let $\mathbf{u}_{0}\in\mathbf{H}$ and $\psi_{0}$ solve (6). Then $\psi_{0}\in H^{1}_{0}(\Omega)$, and there holds a bound
$\lVert\psi_{0}\rVert_{H^{1}_{0}(\Omega)}\leq C\lVert\mathbf{u}_{0}\rVert_{\mathbf{H}}.$
If furthermore, $\mathbf{u}_{0}\in\mathbf{V}$, then $\psi_{0}\in H^{2}_{0}(\Omega)$ and there holds the bound
$\lVert\psi_{0}\rVert_{H^{2}_{0}(\Omega)}\leq C\lVert\mathbf{u}_{0}\rVert_{\mathbf{V}}.$
Before we prove the lemma, note that the result $\psi_{0}\in H^{2}_{0}(\Omega)$ only poses a requirement on $\mathbf{u}_{0}$ and we do not require, e.g., convexity of $\Omega$, which would be otherwise a standard assumption to achieve $H^{2}$ regularity.
Proof.
For the case $\mathbf{u}_{0}\in\mathbf{H}$ there is nothing to prove, as this is directly given by the weak formulation of (6). Hence assume now $\mathbf{u}_{0}\in\mathbf{V}$. Our goal is to show the result by extending the proof of [16, Lemma 2]. As a result of the identity (4), there holds
$\operatorname{curl}(\operatorname{\mathbf{curl}}\psi_{0}-\mathbf{u}_{0})=0.$
By [16, Lemma 3] there exists a unique $p\in H^{1}(\Omega)/\mathbb{R}$ such that $\operatorname{\mathbf{curl}}\psi_{0}-\mathbf{u}_{0}=\nabla p$. Due to the first isomorphism of (10) $\operatorname{\mathbf{curl}}\psi_{0}\in\mathbf{H}$ and by assumption $\mathbf{u}_{0}\in\mathbf{V}\subset\mathbf{H}$, it holds $\nabla p\in\mathbf{H}$. The definition of $\mathbf{H}$ then implies $\Delta p=0$ and $\partial_{\mathbf{n}}p=0$. Hence $p$ is constant and due to $p\in H^{1}(\Omega)/\mathbb{R}$ it actually vanishes. This implies $\operatorname{\mathbf{curl}}\psi_{0}=\mathbf{u}_{0}\in\mathbf{V}\subset H^{1}_{0}(\Omega)^{2}$.
By definition of $\operatorname{\mathbf{curl}}$, this implies
$\partial_{1}\psi_{0},\partial_{2}\psi_{0}\in H^{1}_{0}(\Omega).$
Hence especially $\psi_{0}\in H^{2}(\Omega)$ and $\partial_{\mathbf{n}}\psi_{0}=0$ on $\partial\Omega$. This shows $\psi_{0}\in H^{2}_{0}(\Omega)$. The norm bound can then be obtained using Lemma 1 and Lemma 2, which yields
$\lVert\psi_{0}\rVert_{H^{2}_{0}(\Omega)}\cong\lVert\Delta\psi_{0}\rVert_{L^{2}(\Omega)}=\lVert\operatorname{curl}\mathbf{u}_{0}\rVert_{L^{2}(\Omega)}\cong\lVert\mathbf{u}_{0}\rVert_{\mathbf{V}}.$
2.4 Results for stationary biharmonic equation
Before analyzing the time-dependent problem, we recall some basic results for the biharmonic problem with Dirichlet (clamped) boundary conditions,
$\Delta^{2}w$ $=f\quad$ $\text{in }\Omega,$ (13)
$w=\partial_{n}w$ $=0\quad$ $\text{on }\partial\Omega.$
The weak formulation for (13) reads
$w\in H^{2}_{0}(\Omega)\quad:\quad\left(\Delta w,\Delta\varphi\right)_{\Omega}=\left\langle f,\varphi\right\rangle_{\Omega}\quad\forall\varphi\in H^{2}_{0}(\Omega).$ (14)
The following results can be found, e.g., in [22, Sec. 5.9].
Proposition 5.
Let $\Omega\subset\mathbb{R}^{2}$ be a bounded Lipschitz domain and $f\in H^{-2}(\Omega)$. Then there is a unique weak solution $w\in H^{2}_{0}(\Omega)$ and there exists a constant $C$ such that
$\|w\|_{H^{2}(\Omega)}\leq C\|f\|_{H^{-2}(\Omega)}.$ (15)
If in addition $\Omega$ is convex and $f\in H^{-1}(\Omega)$, then the weak solution $w$ of (13) fulfills $w\in H^{3}(\Omega)\cap H^{2}_{0}(\Omega)$ and there exists a constant $C$ such that
$\|w\|_{H^{3}(\Omega)}\leq C\|f\|_{H^{-1}(\Omega)}.$ (16)
The above result yields, that the operator $\mathcal{A}\colon H^{2}_{0}(\Omega)\to H^{-2}(\Omega)$, defined by
$\left\langle\mathcal{A}w,\varphi\right\rangle_{\Omega}=\left(\Delta w,\Delta\varphi\right)_{\Omega}\qquad\forall\varphi\in H^{2}_{0}(\Omega)$ (17)
is an isomorphism. We immediately obtain the following corollary.
Corollary 6.
Let $g\in H^{1}(\Omega)$ and $w\in H^{2}_{0}(\Omega)$ be the weak solution to the problem
$\left(\Delta w,\Delta\varphi\right)_{\Omega}=\left({\rm NewA}g,{\rm NewA}\varphi\right)_{\Omega}\quad\forall\varphi\in H^{2}_{0}(\Omega).$ (18)
Then there exists a constant $C$ such that
$\|w\|_{H^{2}_{0}(\Omega)}\leq C\|{\rm NewA}g\|_{L^{2}(\Omega)}.$
If $\Omega$ is in addition convex, then $w\in H^{3}(\Omega)\cap H^{2}_{0}(\Omega)$ and
$\|w\|_{H^{3}(\Omega)}\leq C\|{\rm NewA}g\|_{L^{2}(\Omega)}.$ (19)
If $\Omega$ is polygonal, but not necessarily convex, the results of [7, 9] yield, that there exists some $\alpha_{\Omega}>0$, such that for any $0\leq\alpha<\alpha_{\Omega}$, the solution $w$ of (14) for the data $f\in H^{-2+\alpha}(\Omega)$ satisfies a bound
$\|w\|_{H^{2+\alpha}(\Omega)}\leq C\lVert f\rVert_{H^{-2+\alpha}(\Omega)}.$
We consider also an associated eigenvalue problem, the so called buckling eigenvalue problem
$\Delta^{2}w$ $=-\lambda\Delta w$ $\text{in }\Omega,$ (20)
$w=\partial_{n}w$ $=0$ $\text{on }\partial\Omega,$
see, e.g.,[32, 6].
Lemma 7.
There exists a sequence of eigenvalues of (20)
$0<\lambda_{1}\leq\lambda_{2}\leq\dots,\quad\text{with }\lambda_{k}\to\infty,\;\text{for }k\to\infty$
and the corresponding eigenfunctions $w_{k}\in H^{2}_{0}(\Omega)$, which are orthogonal with respect to both $H^{2}_{0}(\Omega)$ and $H^{1}_{0}(\Omega)$ inner products, i.e.
$\left(\Delta w_{k},\Delta w_{j}\right)_{\Omega}=\left(\nabla w_{k},\nabla w_{j}\right)_{\Omega}=0\quad\text{for }k\neq j$
and form basis for both $H^{2}_{0}(\Omega)$ and $H^{1}_{0}(\Omega)$.
Proof.
We consider a bounded linear operator $T\colon H^{1}_{0}(\Omega)\to H^{1}_{0}(\Omega)$ defined as follows:
$T(g)\in H^{2}_{0}(\Omega)\quad:\quad\left(\Delta T(g),\Delta\phi\right)_{\Omega}=\left(\nabla g,\nabla\phi\right)_{\Omega}\quad\text{for all }\phi\in H^{2}_{0}(\Omega).$ (21)
Note, that by Corollary 6 for every $g\in H^{1}_{0}(\Omega)$, $T(g)\in H^{2}_{0}(\Omega)$ is well-defined. Since $H^{2}_{0}(\Omega)$ is compactly embedded into $H^{1}_{0}(\Omega)$ the operator $T\colon H^{1}_{0}(\Omega)\to H^{1}_{0}(\Omega)$ is compact. It is straightforward to check that $T$ is symmetric with respect to the $H^{1}_{0}(\Omega)$ inner product, i.e.
$\left(\nabla T(g),\nabla\phi\right)_{\Omega}=\left(\nabla g,\nabla T(\phi)\right)_{\Omega}\quad\text{for all }g,\phi\in H^{1}_{0}(\Omega),$
and positive definite, since
$\left(\nabla T(g),\nabla g\right)_{\Omega}=\|\Delta T(g)\|^{2}_{L^{2}(\Omega)}\geq 0$
and $\left(\nabla T(g),\nabla g\right)_{\Omega}=0$ implies $T(g)=0$ and thus $g=0$. Thus, $T$ is also injective. We apply spectral theorem for compact normal operators, see, e.g., [5, Theorem 12.12], and obtain the result of the lemma. ∎
Remark 8.
The image of the operator $T\colon H^{1}_{0}(\Omega)\to H^{1}_{0}(\Omega)$ defined in (21) is given as
[…]
By Corollary 6 there holds $\operatorname{Im}(T)=H^{3}(\Omega)\cap H^{2}_{0}(\Omega)$ if $\Omega$ is convex. Since the operator $T$ is injective we can define the inverse
$T^{-1}\colon\operatorname{Im}(T)\to H^{1}_{0}(\Omega),$
which fulfills
$\left(\nabla T^{-1}(w),\nabla\phi\right)_{\Omega}=\left(\Delta w,\Delta\phi\right)_{\Omega}\quad\text{for all }\phi\in H^{2}_{0}(\Omega).$
In the numerical analysis below, we use an operator $A_{h}$ (43), which is a discrete analog of $T^{-1}$.
2.5 Results for time-dependent problems
We call $\psi$ a weak solution of (5) if
[…]
fulfills
$-\langle\partial_{t}\Delta\psi,\varphi\rangle+(\Delta\psi,\Delta\varphi)_{I\times\Omega}$ $=\langle f,\varphi\rangle\quad$ $\text{for all }\;\varphi\in\mathbb{Y}=L^{2}(I;H^{2}_{0}(\Omega)),$ (22)
$\psi(0)$ $=\psi_{0}.$
Here, $\langle\cdot,\cdot\rangle$ denotes the duality product between $L^{2}(I;H^{-2}(\Omega))$ and $L^{2}(I;H^{2}_{0}(\Omega))$.
Before we prove the existence and regularity results for (22), we show two technical results regarding continuity in time, if some Hilbert space regularities are known.
Lemma 9.
Let $\psi\in L^{2}(I;H^{2}_{0}(\Omega))$ satisfy $\partial_{t}\Delta\psi\in L^{2}(I;H^{-2}(\Omega))$. Then (after a possible redefinition on a zero measure subset of $I$) $\psi\in C(\bar{I};H^{1}_{0}(\Omega))$ and there holds a bound
$\lVert\psi\rVert_{C(\bar{I};H^{1}_{0}(\Omega))}\leq C\left(\lVert\psi\rVert_{L^{2}(I;H^{2}_{0}(\Omega))}+\lVert\partial_{t}\Delta\psi\rVert_{L^{2}(I;H^{-2}(\Omega))}\right).$
Proof.
We adapt the proof of [13, Section 5.9, Theorem 3], and for any $\varepsilon>0$ introduce a mollifier $\eta_{\varepsilon}$, which leads to the mollified functions $\psi_{\varepsilon}=\eta_{\varepsilon}*\psi$. For $\varepsilon,\delta>0$ it holds due to the smoothness in time of $\psi_{\varepsilon},\psi_{\delta}$ and $\psi(s)\in H^{2}_{0}(\Omega)$ for a.e. $s\in I$
$\partial_{t}\left(\lVert\nabla\left(\psi_{\varepsilon}(t)-\psi_{\delta}(t)\right)\rVert_{L^{2}(\Omega)}^{2}\right)$ $=2\left(\partial_{t}\nabla(\psi_{\varepsilon}(t)-\psi_{\delta}(t)),\nabla(\psi_{\varepsilon}(t)-\psi_{\delta}(t))\right)_{\Omega}$
$=-2\left(\partial_{t}\Delta(\psi_{\varepsilon}(t)-\psi_{\delta}(t)),(\psi_{\varepsilon}(t)-\psi_{\delta}(t))\right)_{\Omega}$
for all $t\in I$. By the fundamental theorem of calculus, it holds for all $s,t\in I$
$\lVert\nabla\left(\psi_{\varepsilon}(t)-\psi_{\delta}(t)\right)\rVert_{L^{2}(\Omega)}^{2}=\lVert\nabla\left(\psi_{\varepsilon}(s)-\psi_{\delta}(s)\right)\rVert_{L^{2}(\Omega)}^{2}-\int_{s}^{t}2\left(\partial_{t}\Delta(\psi_{\varepsilon}(\tau)-\psi_{\delta}(\tau)),(\psi_{\varepsilon}(\tau)-\psi_{\delta}(\tau))\right)_{\Omega}\,d\tau.$
In the above identity, we fix now some $s\in I$ for which $\psi_{\varepsilon}(s)\to\psi(s)$ in $H^{1}_{0}(\Omega)$. As a consequence, it holds
$\limsup_{\varepsilon,\delta\to 0}\sup_{t\in I}\lVert\nabla\left(\psi_{\varepsilon}(t)-\psi_{\delta}(t)\right)\rVert_{L^{2}(\Omega)}^{2}\leq 2\lim_{\varepsilon,\delta\to 0}\lVert\partial_{t}\Delta(\psi_{\varepsilon}-\psi_{\delta})\rVert_{L^{2}(I;H^{-2}(\Omega))}\lVert\psi_{\varepsilon}-\psi_{\delta}\rVert_{L^{2}(I;H^{2}_{0}(\Omega))}=0.$
Hence as in the proof of the cited reference, this yields that (after a possible redefinition on a set of measure zero) it holds
$\nabla\psi\in C(\bar{I};L^{2}(\Omega)^{2}).$ (23)
Now that we have fixed a representation of $\psi$ with $\nabla\psi$ being continuous in time, we have to take special care with further redefinitions on zero sets. To conclude the proof, note that $\psi\in L^{2}(I;H^{2}_{0}(\Omega))$ especially implies that $\psi(t)\in H^{1}_{0}(\Omega)$ for a.e. $t\in I$. Hence, for every $t^{*}\in\bar{I}$, there exists a sequence $\{t_{n}\}\subset I$ with $t_{n}\to t^{*}$, $\psi(t_{n})\in H^{1}_{0}(\Omega)$ for all $n$ and by (23)
$\lVert\nabla\psi(t_{n})-\nabla\psi(t^{*})\rVert_{L^{2}(\Omega)}\to 0,\ n\to\infty.$ (24)
Especially $\{\psi(t_{n})\}$ is a Cauchy sequence in $H^{1}_{0}(\Omega)$ and thus $\psi(t_{n})\to\psi^{*}\in H^{1}_{0}(\Omega)$ for some $\psi^{*}\in H^{1}_{0}(\Omega)$. From (24), we deduce $\nabla\psi(t^{*})=\nabla\psi^{*}$, and especially that $\psi^{*}$ is independent of the explicit choice of the sequence $\{t_{n}\}$. Thus, we can redefine $\psi(t^{*}):=\psi^{*}$, without breaking the property (23). ∎
Lemma 10.
Let $\psi\in L^{2}(I;H^{2}_{0}(\Omega))$ satisfy $\partial_{t}\psi\in L^{2}(I;H^{1}_{0}(\Omega))$ and $\mathcal{A}\psi\in L^{2}(I;H^{-1}(\Omega))$. Then (after a possible redefinition on a measure zero subset of $I$) $\psi\in C(\bar{I};H^{2}_{0}(\Omega))$ and there holds a bound
$\lVert\psi\rVert_{C(\bar{I};H^{2}_{0}(\Omega))}\leq C\left(\lVert\Delta^{2}\psi\rVert_{L^{2}(I;H^{-1}(\Omega))}+\lVert\partial_{t}\psi\rVert_{L^{2}(I;H^{1}_{0}(\Omega))}\right).$
Proof.
Similarly to the proof of Lemma 9, we consider the mollified functions $\psi_{\varepsilon}=\psi*\eta_{\varepsilon}$. As $\partial_{t}(\psi*\eta_{\varepsilon})=\psi*(\partial_{t}\eta_{\varepsilon})$, and $\psi(t)\in H^{2}_{0}(\Omega)$ for a.e. $t\in I$, it holds $\partial_{t}(\psi*\eta_{\varepsilon})\in H^{2}_{0}(\Omega)$ for all $t\in I$, and thus we obtain via integration by parts:
$\partial_{t}\left(\lVert\Delta\left(\psi_{\varepsilon}(t)-\psi_{\delta}(t)\right)\rVert_{L^{2}(\Omega)}^{2}\right)$ $=2\left(\partial_{t}\Delta(\psi_{\varepsilon}(t)-\psi_{\delta}(t)),\Delta(\psi_{\varepsilon}(t)-\psi_{\delta}(t))\right)_{\Omega}$
$=2\left(\partial_{t}(\psi_{\varepsilon}(t)-\psi_{\delta}(t)),\mathcal{A}(\psi_{\varepsilon}(t)-\psi_{\delta}(t))\right)_{\Omega}.$
By the fundamental theorem of calculus, it holds for all $s,t\in I$
$\lVert\Delta\left(\psi_{\varepsilon}(t)-\psi_{\delta}(t)\right)\rVert_{L^{2}(\Omega)}^{2}=\lVert\Delta\left(\psi_{\varepsilon}(s)-\psi_{\delta}(s)\right)\rVert_{L^{2}(\Omega)}^{2}+\int_{s}^{t}2\left(\partial_{t}(\psi_{\varepsilon}(\tau)-\psi_{\delta}(\tau)),\mathcal{A}(\psi_{\varepsilon}(\tau)-\psi_{\delta}(\tau))\right)_{\Omega}\,d\tau.$
Hence, fixing any $s\in I$ with $\Delta\psi_{\varepsilon}(s)\to\Delta\psi(s)$ in $L^{2}(\Omega)$:
$\limsup_{\varepsilon,\delta\to 0}\sup_{t\in I}\lVert\Delta\left(\psi_{\varepsilon}(t)-\psi_{\delta}(t)\right)\rVert_{L^{2}(\Omega)}^{2}\leq\lim_{\varepsilon,\delta\to 0}\lVert\partial_{t}(\psi_{\varepsilon}-\psi_{\delta})\rVert_{L^{2}(I;H^{1}_{0}(\Omega))}\lVert\mathcal{A}(\psi_{\varepsilon}-\psi_{\delta})\rVert_{L^{2}(I;H^{-1}(\Omega))}=0.$
Thus, we obtain as before, that
$\Delta\psi\in C(\bar{I};L^{2}(\Omega)).$
As in the proof of Lemma 9, using this result, and the fact that $H^{2}_{0}(\Omega)\subset H^{2}(\Omega)$ is closed, we deduce that in fact $\psi\in C(\bar{I};H^{2}_{0}(\Omega))$. ∎
For (22) we have the following existence and stability result.
Theorem 11 (Existence, uniqueness and stability).
Let $f\in L^{2}(I;H^{-2}(\Omega))$ and $\psi_{0}\in H^{1}_{0}(\Omega)$. Then there exists a unique solution $\psi\in\mathbb{X}$ to (22). Moreover, $\psi\in C(\bar{I};H^{1}_{0}(\Omega))$ and the following estimate holds
$\|\partial_{t}\Delta\psi\|_{L^{2}(I;H^{-2}(\Omega))}+\|\psi\|_{L^{2}(I;H^{2}_{0}(\Omega))}+\|\psi\|_{C(\bar{I};H^{1}_{0}(\Omega))}\leq C\left(\|f\|_{L^{2}(I;H^{-2}(\Omega))}+\|{\rm NewA}\psi_{0}\|_{L^{2}(\Omega)}\right).$
Proof.
The proof uses a Galerkin procedure similar to the proof of [13, Section 7.1.2, Theorem 3] in the case of the heat equation. We use the eigenfunctions $w_{k}\in H^{2}_{0}(\Omega)$ from Lemma 7 and normalize them such that $\|\nabla w_{k}\|_{L^{2}(\Omega)}=1$. For $n\in\mathbb{N}$ we consider the space
[…]
and a Galerkin approximation to (22) given as $\psi_{n}\colon\bar{I}\to V_{n}$ and fulfilling
$\left(\nabla\partial_{t}\psi_{n}(t),\nabla v_{n}\right)_{\Omega}+\left(\Delta\psi_{n}(t),\Delta v_{n}\right)_{\Omega}$ $=\langle f(t),v_{n}\rangle_{H^{-2}(\Omega),H^{2}_{0}(\Omega)}\quad$ $\text{for all }\;v_{n}\in V_{n}\text{ and a.a. }t\in I$ (25)
$\left(\nabla\psi_{n}(0),\nabla v_{n}\right)_{\Omega}$ $=\left(\nabla\psi_{0},\nabla v_{n}\right)_{\Omega}\quad$ $\text{for all }\;v_{n}\in V_{n}.$
The existence and uniqueness of a solution of this systems of ODEs follows directly. The initial condition in (25), can compactly be written as
$\psi_{n}(0)=P_{n}\psi_{0},$ (26)
where $P_{n}\colon H^{1}_{0}(\Omega)\to V_{n}$ is the orthogonal projection defined by
$\left(\nabla P_{n}v,\nabla v_{n}\right)_{\Omega}=\left(\nabla v,\nabla v_{n}\right)_{\Omega}\qquad\text{for all $v_{n}\in V_{n}$}.$ (27)
It is straightforward to check, that this projection is stable in $H^{1}_{0}(\Omega)$, i.e., there holds
$\|P_{n}v\|_{H^{1}_{0}(\Omega)}\leq\|v\|_{H^{1}_{0}(\Omega)}.$ (28)
Testing (25) with $v_{n}=\psi_{n}(t)$ we obtain
$\frac{1}{2}\frac{d}{dt}\|{\rm NewA}\psi_{n}(t)\|^{2}_{L^{2}(\Omega)}+\|\Delta\psi_{n}(t)\|^{2}_{L^{2}(\Omega)}=\langle f(t),\psi_{n}(t)\rangle_{H^{-2}(\Omega),H^{2}_{0}(\Omega)}\leq\|f(t)\|_{H^{-2}(\Omega)}\|\psi_{n}(t)\|_{H^{2}(\Omega)}\quad\text{a.e. $t\in I$}.$
Integrating in time over $(0,t)$ for arbitrary $t\in I$, and using the fact that due to Lemma 1 $\|\psi_{n}(t)\|_{H^{2}_{0}(\Omega)}=\|\Delta\psi_{n}(t)\|_{L^{2}(\Omega)}$ we obtain after an application of the Young’s inequality
$\frac{1}{2}\|{\rm NewA}\psi_{n}(t)\|^{2}_{L^{2}(\Omega)}+\|\psi_{n}\|^{2}_{L^{2}((0,t);H^{2}_{0}(\Omega))}\leq\frac{1}{2}\|{\rm NewA}\psi_{n}(0)\|^{2}_{L^{2}(\Omega)}+\frac{1}{2}\|\psi_{n}\|^{2}_{L^{2}((0,t);H^{2}_{0}(\Omega))}+\frac{1}{2}\|f\|^{2}_{L^{2}((0,t);H^{-2}(\Omega))}.$
Using the fact that, due to (28) $\|\nabla\psi_{n}(0)\|^{2}_{L^{2}(\Omega)}\leq\|\nabla\psi_{0}\|^{2}_{L^{2}(\Omega)}$, and $t\in I$ was arbitrary, we get
$\|{\rm NewA}\psi_{n}\|^{2}_{L^{\infty}(I;L^{2}(\Omega))}+\|\psi_{n}\|^{2}_{L^{2}(I;H^{2}_{0}(\Omega))}\leq\|{\rm NewA}\psi_{0}\|^{2}_{L^{2}(\Omega)}+\|f\|^{2}_{L^{2}(I;H^{-2}(\Omega))}.$ (29)
In the next step we show the boundedness of $\|\partial_{t}\Delta\psi_{n}\|_{L^{2}(I;H^{-2}(\Omega))}$. By the orthonormality in $H^{1}_{0}(\Omega)$ of the basis we get for every $v\in H^{1}_{0}(\Omega)$
$v=\sum_{k=1}^{\infty}\left(\nabla v,\nabla w_{k}\right)_{\Omega}w_{k}\quad\text{and}\quad P_{n}v=\sum_{k=1}^{n}\left(\nabla v,\nabla w_{k}\right)_{\Omega}w_{k}.$
Moreover, if $v\in H^{2}_{0}(\Omega)$ and $v_{n}\in V_{n}$, then by orthogonality of $\{w_{k}\}$ in $H^{2}_{0}(\Omega)$:
$\left(\Delta v_{n},\Delta v\right)_{\Omega}=\left(\Delta v_{n},\sum_{k=1}^{\infty}\Delta w_{k}\left({\rm NewA}v,{\rm NewA}w_{k}\right)_{\Omega}\right)_{\Omega}=\left(\Delta v_{n},\sum_{k=1}^{n}\Delta w_{k}\left({\rm NewA}v,{\rm NewA}w_{k}\right)_{\Omega}\right)_{\Omega}=\left(\Delta v_{n},\Delta P_{n}v\right)_{\Omega}.$ (30)
By the same arguments, we obtain for any $v\in H^{2}_{0}(\Omega)$
$\|P_{n}v\|_{H^{2}_{0}(\Omega)}^{2}=\|\Delta P_{n}v\|^{2}_{L^{2}(\Omega)}=\sum_{k=1}^{n}\left(\nabla v,\nabla w_{k}\right)_{\Omega}^{2}\|\Delta w_{k}\|^{2}_{L^{2}(\Omega)}\leq\sum_{k=1}^{\infty}\left(\nabla v,\nabla w_{k}\right)_{\Omega}^{2}\|\Delta w_{k}\|^{2}_{L^{2}(\Omega)}=\|\Delta v\|^{2}_{L^{2}(\Omega)}.$
Let $v\in H^{2}_{0}(\Omega)$ be arbitrary. As $\psi_{n}(t)\in V_{n}$ a.e. in $I$, there holds $\partial_{t}\psi_{n}(t)\in V_{n}$ a.e. in $I$, and thus for a.e. $t\in I$, we get
$(\partial_{t}\Delta\psi_{n}(t),v)_{\Omega}$ $=-(\nabla\partial_{t}\psi_{n}(t),\nabla v)_{\Omega}=-(\nabla\partial_{t}\psi_{n}(t),\nabla P_{n}v)_{\Omega}$
$=(\Delta\psi_{n}(t),\Delta P_{n}v)_{\Omega}-\langle f(t),P_{n}v\rangle_{H^{-2}(\Omega),H^{2}_{0}(\Omega)}$
$\leq\|\Delta\psi_{n}(t)\|_{L^{2}(\Omega)}\|\Delta P_{n}v\|_{L^{2}(\Omega)}+\|f(t)\|_{H^{-2}(\Omega)}\|P_{n}v\|_{H^{2}_{0}(\Omega)}$
$\leq\left(\|\psi_{n}(t)\|_{H^{2}_{0}(\Omega)}+\|f(t)\|_{H^{-2}(\Omega)}\right)\|v\|_{H^{2}_{0}(\Omega)}.$
Thus for a.e. $t\in I$, there holds
$\|\partial_{t}\Delta\psi_{n}(t)\|_{H^{-2}(\Omega)}\leq\|\psi_{n}(t)\|_{H^{2}_{0}(\Omega)}+\|f(t)\|_{H^{-2}(\Omega)}.$
Squaring, integrating w.r.t $t$ and using (29), we get
$\|\partial_{t}\Delta\psi_{n}\|_{L^{2}(I;H^{-2}(\Omega))}\leq C\left(\|{\rm NewA}\psi_{0}\|_{L^{2}(\Omega)}+\|f\|_{L^{2}(I;H^{-2}(\Omega))}\right).$ (31)
From these estimates we obtain existence of a subsequence of $\psi_{n}$ (denoted again by $\psi_{n}$) and some $\psi\in\mathbb{X}$ such that
$\psi_{n}\rightharpoonup\psi\text{ in }L^{2}(I;H^{2}_{0}(\Omega))\;\text{ and }\;\partial_{t}\Delta\psi_{n}\rightharpoonup\partial_{t}\Delta\psi\text{ in }L^{2}(I;H^{-2}(\Omega)).$ (32)
To conclude the proof, let $\varphi^{N}\in C^{1}(\bar{I};H^{2}_{0}(\Omega))$ be of the form $\varphi^{N}=\sum_{j=1}^{N}\zeta_{j}(t)w_{j}$ where $N\in\mathbb{N}$ is fixed and $\zeta_{i}(t)\in C^{1}(\bar{I})$. Then, testing the first equation of (25) for $n>N$ with $\varphi^{N}(t)$ and integrating over $I$ yields
$\left\langle\partial_{t}\Delta\psi_{n},\varphi^{N}\right\rangle_{I\times\Omega}+\left(\Delta\psi_{n},\Delta\varphi^{N}\right)_{I\times\Omega}=\langle f,\varphi^{N}\rangle.$
Using the weak convergences (32) thus yields
$\left\langle\partial_{t}\Delta\psi,\varphi^{N}\right\rangle_{I\times\Omega}+\left(\Delta\psi,\Delta\varphi^{N}\right)_{I\times\Omega}=\langle f,\varphi^{N}\rangle.$
As functions of the form of $\varphi^{N}$ are dense in $\mathbb{Y}=L^{2}(I;H^{2}_{0}(\Omega))$, we have just shown the first equation of (22). To that the initial condition is satisfied, let $\varphi\in C^{1}(\bar{I};H^{2}_{0}(\Omega))$ be arbitrary with $\varphi(T)=0$. Testing the first equation of (22) with $\varphi$, yields after integration by parts in time and space, that
$\left\langle\nabla\psi,\partial_{t}\nabla\varphi\right\rangle_{I\times\Omega}+\left(\Delta\psi,\Delta\varphi\right)_{I\times\Omega}=\langle f,\varphi\rangle+\left(\nabla\psi(0),\nabla\varphi(0)\right)_{\Omega}.$
Moreover, using integration by parts in time, $\varphi(T)=0$, the definition of $P_{n}$ (27), as well as (30) applied to every $t\in I$, we have
$-\left\langle\nabla\psi_{n},\partial_{t}\nabla\varphi\right\rangle_{I\times\Omega}+\left(\Delta\psi_{n},\Delta\varphi\right)_{I\times\Omega}$ $=\left\langle\partial_{t}\nabla\psi_{n},\nabla\varphi\right\rangle_{I\times\Omega}+\left(\Delta\psi_{n},\Delta\varphi\right)_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla\varphi(0)\right)_{\Omega}$
$=\left\langle\partial_{t}\nabla\psi_{n},\nabla P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\Delta\psi_{n},\Delta P_{n}\varphi\right)_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla P_{n}\varphi(0)\right)_{\Omega}$
$=\left\langle f,P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla P_{n}\varphi(0)\right)_{\Omega}$
$=\left\langle f,P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\nabla\psi_{n}(0),\nabla\varphi(0)\right)_{\Omega}.$
Hence, subtracting the two equations, and taking the limit yields due to the weak convergence (32) and (28)
$0$ $=\lim_{n\to\infty}\left[\left\langle f,\varphi-P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\nabla(\psi(0)-\psi_{n}(0)),\nabla\varphi(0)\right)_{\Omega}\right]$
$=\lim_{n\to\infty}\left[\left\langle f,\varphi-P_{n}\varphi\right\rangle_{I\times\Omega}+\left(\nabla(\psi(0)-P_{n}\psi_{0}),\nabla\varphi(0)\right)_{\Omega}\right].$
Using the convergence $P_{n}\varphi(t)\to\varphi(t)$ in $H^{2}_{0}(\Omega)$ for all $t\in I$, together with dominated convergence, as well as $P_{n}\psi_{0}\to\psi_{0}$ in $H^{1}_{0}(\Omega)$ yields
$\left(\nabla(\psi(0)-\psi_{0}),\nabla\varphi(0)\right)_{\Omega}=0.$
As $\varphi(0)$ can be arbitrary, we have concluded with the proof of $\psi(0)=\psi_{0}$. From the estimates (29),(31), we know that $\psi$ satisfies the bounds
$\|\partial_{t}\Delta\psi\|_{L^{2}(I;H^{-2}(\Omega))}+\|\psi\|_{L^{2}(I;H^{2}_{0}(\Omega))}\leq C\left(\|f\|_{L^{2}(I;H^{-2}(\Omega))}+\|{\rm NewA}\psi_{0}\|_{L^{2}(\Omega)}\right).$
The continuity in time is then a direct consequence of Lemma 9. Choosing $\psi_{0}=0$ and $f=0$ directly implies uniqueness. ∎
Remark 12.
The stability estimate of Theorem 11 bounds the term $\|\partial_{t}\Delta\psi\|_{L^{2}(I;H^{-2}(\Omega))}$ by norms of the data. From this it is however not clear, whether $\partial_{t}\psi\in L^{2}(I\times\Omega)$ holds. In fact such a result seems doubtful, as by a duality argument this would relate to $\Delta^{-1}\colon L^{2}(\Omega)\to H^{2}_{0}(\Omega)$ which does not hold in general even on convex domains.
Under some additional conditions on data we can show higher regularity result.
Theorem 13.
Let $f\in L^{2}(I;H^{-1}(\Omega))$ and $\psi_{0}\in H^{2}_{0}(\Omega)$. Then the solution to (22) additionally has the following regularities: $\partial_{t}\psi\in L^{2}(I;H^{1}_{0}(\Omega))$, $\mathcal{A}\psi\in L^{2}(I;H^{-1}(\Omega))$, and $\psi\in C(\bar{I};H^{2}_{0}(\Omega))$. The weak form can be written as
$\left(\partial_{t}{\rm NewA}\psi,{\rm NewA}\varphi\right)_{I\times\Omega}+\left(\Delta\psi,\Delta\varphi\right)_{I\times\Omega}$ $=\langle f,\varphi\rangle\quad$ $\text{for all }\;\varphi\in\mathbb{Y},$ (33)
$\psi(0)$ $=\psi_{0},$
and there holds
$\|{\rm NewA}\partial_{t}\psi\|_{L^{2}(I;L^{2}(\Omega))}+\|\mathcal{A}\psi\|_{L^{2}(I;H^{-1}(\Omega))}+\|\psi\|_{C(\bar{I};H^{2}_{0}(\Omega))}\leq C\left(\|f\|_{L^{2}(I;H^{-1}(\Omega))}+\|\psi_{0}\|_{H^{2}_{0}(\Omega)}\right).$
If $\Omega$ is in addition convex, then $\psi\in L^{2}(I;H^{3}(\Omega)\cap H^{2}_{0}(\Omega))$ and there holds
$\|\psi\|_{L^{2}(I;H^{3}(\Omega))}\leq C\left(\|f\|_{L^{2}(I;H^{-1}(\Omega))}+\|\psi_{0}\|_{H^{2}_{0}(\Omega)}\right).$
Proof.
We use the Galerkin approximation from the proof of Theorem 11 and take $v_{n}=\partial_{t}\psi_{n}(t)\in V_{n}$. This results in
$\|{\rm NewA}\partial_{t}\psi_{n}(t)\|^{2}_{L^{2}(\Omega)}+\frac{1}{2}\frac{d}{dt}\|\Delta\psi_{n}(t)\|^{2}_{L^{2}(\Omega)}=\langle f(t),\partial_{t}\psi_{n}(t)\rangle_{H^{-1}(\Omega),H^{1}_{0}(\Omega)}\leq\|f(t)\|_{H^{-1}(\Omega)}\|\nabla\partial_{t}\psi_{n}(t)\|_{L^{2}(\Omega)}.$
Integrating over $I$ and using the Young’s inequality, we have
$\|{\rm NewA}\partial_{t}\psi_{n}\|^{2}_{I\times\Omega}+\frac{1}{2}\|\Delta\psi_{n}(T)\|^{2}_{L^{2}(\Omega)}\leq\frac{1}{2}\|\Delta\psi_{n}(0)\|^{2}_{L^{2}(\Omega)}+\frac{1}{2}\|\nabla\partial_{t}\psi_{n}\|^{2}_{I\times\Omega}+\frac{1}{2}\|f\|^{2}_{L^{2}(I;H^{-1}(\Omega))}.$ (34)
As $\psi_{0}\in H^{2}_{0}(\Omega)$, due to (26) and (30), there holds
$\|\Delta\psi_{n}(0)\|^{2}_{L^{2}(\Omega)}=\|\Delta P_{n}\psi_{0}\|^{2}_{L^{2}(\Omega)}\leq\|\Delta\psi_{0}\|^{2}_{L^{2}(\Omega)}=\|\psi_{0}\|^{2}_{H^{2}_{0}(\Omega)}.$
Hence, after canceling terms in (34), multiplying by 2 and taking the root, we obtain
$\|{\rm NewA}\partial_{t}\psi_{n}\|_{I\times\Omega}\leq\|\psi_{0}\|_{H^{2}_{0}(\Omega)}+\|f\|_{L^{2}(I;H^{-1}(\Omega))}.$ (35)
Moreover, using integration by parts, (30) and the definition of the Galerkin approximation (25), there holds for any $v\in H^{2}_{0}(\Omega)$ and a.e. $t\in I$
$\left(\mathcal{A}\psi_{n}(t),v\right)_{\Omega}$ $=\left(\Delta\psi_{n}(t),\Delta v\right)_{\Omega}=\left(\Delta\psi_{n}(t),\Delta P_{n}v\right)_{\Omega}=-\left(\partial_{t}{\rm NewA}\psi_{n}(t),{\rm NewA}P_{n}v\right)_{\Omega}+\langle f(t),P_{n}v\rangle.$
By density of $H^{2}_{0}(\Omega)$ in $H^{1}_{0}(\Omega)$, for all $v\in H^{1}_{0}(\Omega)$ we thus have
$\left(\mathcal{A}\psi_{n}(t),v\right)_{\Omega}=-\left(\partial_{t}{\rm NewA}\psi_{n}(t),{\rm NewA}P_{n}v\right)_{\Omega}+\langle f(t),P_{n}v\rangle.$
Hence, using the stability of $P_{n}$, (28), we obtain the bound
$\left(\mathcal{A}\psi_{n}(t),v\right)_{\Omega}$ $\leq C\left(\lVert\partial_{t}{\rm NewA}\psi_{n}(t)\rVert_{L^{2}(\Omega)}+\lVert f(t)\rVert_{H^{-1}(\Omega)}\right)\lVert P_{n}v\rVert_{H^{1}_{0}(\Omega)}$
$\leq C\left(\lVert\partial_{t}{\rm NewA}\psi_{n}(t)\rVert_{L^{2}(\Omega)}+\lVert f(t)\rVert_{H^{-1}(\Omega)}\right)\lVert v\rVert_{H^{1}_{0}(\Omega)},$
which leads to the estimate
$\lVert\mathcal{A}\psi_{n}(t)\rVert_{H^{-1}(\Omega)}\leq C\left(\lVert\partial_{t}{\rm NewA}\psi_{n}(t)\rVert_{L^{2}(\Omega)}+\lVert f(t)\rVert_{H^{-1}(\Omega)}\right),$
and thus after squaring and integrating in time, and using (35)
$\lVert\mathcal{A}\psi_{n}\rVert_{L^{2}(I;H^{-1}(\Omega))}\leq C\left(\|\psi_{0}\|_{H^{2}(\Omega)}+\|f\|_{L^{2}(I;H^{-1}(\Omega))}\right).$ (36)
As in the proof of Theorem 11, we obtain $\mathcal{A}\psi_{n}\rightharpoonup\mathcal{A}\psi$ in $L^{2}(I;L^{2}(\Omega))$ and ${\rm NewA}\partial_{t}\psi_{n}\rightharpoonup{\rm NewA}\partial_{t}\psi$ in $L^{2}(I;H^{-1}(\Omega)^{2})$ where $\psi$ is the solution to (22). As balls in the norms of $L^{2}(I;L^{2}(\Omega))$ and $L^{2}(I;H^{-1}(\Omega)^{2})$ are weakly closed, the bounds (35) and (36) derived for $\psi_{n}$ also hold for $\psi$. Using $\partial_{t}\psi\in L^{2}(I;H^{1}_{0}(\Omega))$ and $\mathcal{A}\psi\in L^{2}(I;H^{-1}(\Omega))$ together with Lemma 10 yields $\psi\in C(\bar{I};H^{2}_{0}(\Omega))$. If $\Omega$ is convex, due to $\mathcal{A}\psi\in L^{2}(I;H^{-1}(\Omega))$, we may apply the $H^{3}$ regularity from Proposition 5 for a.e. $t\in I$. This yields $\psi\in L^{2}(I;H^{3}(\Omega)\cap H^{2}_{0}(\Omega))$ together with the corresponding estimate. ∎
Remark 14.
In Theorems 11 and 13, we have stated the stability results in terms of the data for the stream-function formulation. If the data $f,\psi_{0}$ are not directly given, but rather have to be deduced from the data for the corresponding Stokes problem, i.e. it holds $f=\operatorname{curl}(\mathbf{g})$ and $\psi_{0}$ solves (6), we have the following relations. If $\mathbf{g}\in L^{2}(I;\mathbf{V}^{*})$ and $\mathbf{u}_{0}\in\mathbf{H}$, then
$\|\partial_{t}\Delta\psi\|_{L^{2}(I;H^{-2}(\Omega))}+\|\psi\|_{L^{2}(I;H^{2}_{0}(\Omega))}+\|\psi\|_{C(\bar{I};H^{1}(\Omega))}\leq C\left(\|\mathbf{g}\|_{L^{2}(I;\mathbf{V}^{*})}+\|\mathbf{u}_{0}\|_{L^{2}(\Omega)}\right).$
If moreover $\mathbf{g}\in L^{2}(I;\mathbf{H})$ and $\mathbf{u}_{0}\in\mathbf{V}$, then
$\|{\rm NewA}\partial_{t}\psi\|_{L^{2}(I;L^{2}(\Omega))}+\|\mathcal{A}\psi\|_{L^{2}(I;H^{-1}(\Omega))}+\|\psi\|_{C(\bar{I};H^{2}_{0}(\Omega))}\leq C\left(\|\mathbf{g}\|_{L^{2}(I;\mathbf{H})}+\|\mathbf{u}_{0}\|_{\mathbf{V}}\right).$
These are direct consequences of Lemma 4 and (11).
2.6 Interpolated Regularity
The formulations of Theorems 11 and 13 were chosen to explicitly highlight the best regulsrity of $\psi$ that we require, but obscures the derivation of a corresponding intermediate result. Before we state the next corollary, let us note, that the stability estimate of Theorem 11 can equivalently be written as: If $f\in L^{2}(I;H^{-2}(\Omega))$ and $\psi_{0}\in H^{1}_{0}(\Omega)$, then
$\|\Delta\partial_{t}\psi\|_{L^{2}(I;H^{-2}(\Omega))}+\|\mathcal{A}\psi\|_{L^{2}(I;H^{-2}(\Omega))}\leq C\left(\|\psi_{0}\|_{H^{1}_{0}(\Omega)}+\|f\|_{L^{2}(I;H^{-2}(\Omega))}\right).$ (37)
Accordingly, Theorem 13 states that, if $f\in L^{2}(I;H^{-1}(\Omega))$ and $\psi_{0}\in H^{2}_{0}(\Omega)$, then
$\|\Delta\partial_{t}\psi\|_{L^{2}(I;H^{-1}(\Omega))}+\|\mathcal{A}\psi\|_{L^{2}(I;H^{-1}(\Omega))}\leq C\left(\|\psi_{0}\|_{H^{2}_{0}(\Omega)}+\|f\|_{L^{2}(I;H^{-1}(\Omega))}\right),$ (38)
where in both statements we use the notation $\mathcal{A}$ for the biharmonic operator as introduced in (17). We then immediately obtain:
Corollary 15.
Let $\theta\in[0,1]$, $f\in L^{2}(I;H^{-2+\theta}(\Omega))$ and $\psi_{0}\in H^{1+\theta}_{0}(\Omega)$. Then the solution to (22) satisfies the estimate
$\|\Delta\partial_{t}\psi\|_{L^{2}(I;H^{-2+\theta}(\Omega))}+\|\mathcal{A}\psi\|_{L^{2}(I;H^{-2+\theta}(\Omega))}\leq C\left(\|\psi_{0}\|_{H^{1+\theta}_{0}(\Omega)}+\|f\|_{L^{2}(I;H^{-2+\theta}(\Omega))}\right).$ (39)
Proof.
The result is a direct consequence of (37) and (38) together with the complex interpolation. ∎
3 Fully discrete approximation
In this section we consider the discrete version of the operators presented in the previous section and introduce fully discrete Galerkin solution.
3.1 Spatial discretization
Let $\{\mathcal{T}_{h}\}$ be a family of triangulations of $\bar{\Omega}$, consisting of closed simplices, where we denote by $h$ the maximum mesh-size. Let $V_{h}\subset H^{1}_{0}(\Omega)$ be the finite dimensional space. We denote by $\Pi_{h}\colon H^{1}_{0}(\Omega)\to V_{h}$ the $H^{1}_{0}$-projection, defined by
$\left(\nabla\Pi_{h}w,\nabla v_{h}\right)_{\Omega}=\left(\nabla w,\nabla v_{h}\right)_{\Omega}\qquad\forall v_{h}\in V_{h}.$ (40)
By definition, $\Pi_{h}$ is stable in $H^{1}_{0}(\Omega)$, i.e. there holds
$\|\Pi_{h}w\|_{H^{1}_{0}(\Omega)}\leq\|w\|_{H^{1}_{0}(\Omega)}.$ (41)
To allow for various space discretizations, we introduce a symmetric discrete elliptic bilinear form
$a_{h}(\cdot,\cdot):V_{h}\times V_{h}\to\mathbb{R},$
where at this point we do not fix a special choice of discretization in space. As $a_{h}(\cdot,\cdot)$ was defined to be symmetric and elliptic, i.e., there holds
$a_{h}(\varphi_{h},\varphi_{h})\geq 0\quad\forall\varphi_{h}\in V_{h}\quad\text{and}\quad a_{h}(\varphi_{h},\varphi_{h})=0\quad\Rightarrow\quad\varphi_{h}=0,$
the bilinear form $a_{h}(\cdot,\cdot)$ defines a norm on $V_{h}$
[…]
In addition we also define an operator $A_{h}\colon V_{h}\to V_{h}$, by
$a_{h}(\phi_{h},\varphi_{h})=({\rm NewA}A_{h}\phi_{h},{\rm NewA}\varphi_{h})_{\Omega}\quad\forall\phi_{h},\varphi_{h}\in V_{h}.$ (43)
Notice, since the bilinear form $a_{h}(\cdot,\cdot)$ is symmetric, the operator $A_{h}$ is self-adjoint w.r.t the $H^{1}_{0}(\Omega)$ inner product, since
$({\rm NewA}A_{h}\phi_{h},{\rm NewA}\varphi_{h})_{\Omega}=a_{h}(\phi_{h},\varphi_{h})=a_{h}(\varphi_{h},\phi_{h})=({\rm NewA}A_{h}\varphi_{h},{\rm NewA}\phi_{h})_{\Omega}=({\rm NewA}\phi_{h},{\rm NewA}A_{h}\varphi_{h})_{\Omega}.$
We can establish the following stability result
Lemma 16.
Let $g\in H^{1}_{0}(\Omega)$ and $w_{h}\in V_{h}$ be the finite element solution
$a_{h}(w_{h},\varphi_{h})=({\rm NewA}g,{\rm NewA}\varphi_{h})_{\Omega}\quad\forall\varphi_{h}\in V_{h}.$
Then,
$\|{\rm NewA}A_{h}w_{h}\|_{L^{2}(\Omega)}\leq\|{\rm NewA}g\|_{L^{2}(\Omega)}.$
Proof.
By the definition of $A_{h}$ and the equation for $w_{h}$,
$\|{\rm NewA}A_{h}w_{h}\|_{\Omega}^{2}=({\rm NewA}A_{h}w_{h},{\rm NewA}A_{h}w_{h})_{\Omega}=a_{h}(w_{h},A_{h}w_{h})=({\rm NewA}g,{\rm NewA}A_{h}w_{h})_{\Omega}\leq\|{\rm NewA}g\|_{\Omega}\|{\rm NewA}A_{h}w_{h}\|_{\Omega}.$
Canceling, we obtain the result ∎
We wish our analysis to be applicable to a broad range of discretizations, for instance to the $C^{0}$-interior penalty method (cf. [10]), in which case $a_{h}(\psi,\cdot)$ is not well defined for general $\psi\in H^{2}_{0}(\Omega)$. Vice versa, not for all data $f\in H^{-2}(\Omega)$ the right hand side can be tested with $\phi_{h}\in V_{h}$. Hence, throughout the paper, we work under the following assumptions, the first ensuring wellposedness of the discrete problems, and the second guaranteeing an elliptic Galerkin orthogonality.
Assumption 1.
There exists $\theta\in(0,1]$, such that $V_{h}\subset H^{1+\theta}_{0}(\Omega)$.
Assumption 2.
Let $\theta$ and $V_{h}$ be as in 1. For all $\psi\in X:=\mathcal{A}^{-1}(H^{-1-\theta}(\Omega))$, $\varphi_{h}\in V_{h}$, the bilinear form $a_{h}(\cdot,\cdot)$ can be extended to
$a_{h}(\cdot,\cdot):X+V_{h}\times V_{h}\to\mathbb{R},$
and there holds
$\left\langle\mathcal{A}\psi,\varphi_{h}\right\rangle_{\Omega}=a_{h}(\psi,\varphi_{h}),$
where $\mathcal{A}$, is the biharmonic operator introduced in (17).
Remark 17.
Although technical in nature, these assumptions are not overly restrictive, mainly posing a regularity assumption on the solution $\psi$ of (13). In [10, Eq. 1.4] the authors show that on any polygonal domain, for the choice $\theta=\frac{1}{2}-\varepsilon$, where $\varepsilon>0$ can be taken arbitrarily small, there holds $X=H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}+\varepsilon}(\Omega)$. For this choice of $\theta$, taking $V_{h}$ as a standard $C^{0}$-finite element space trivially satisfies 1. In [10, Lemma 5] it is then shown, that if $a_{h}(\cdot,\cdot)$ is chosen as the interior penalty bilinear form then 2 holds. On the other hand, for $H^{2}_{0}(\Omega)\supset V_{h}$ the $C^{1}$-finite element space, together with $a_{h}(\cdot,\cdot)=\left(\Delta\cdot,\Delta\cdot\right)_{\Omega}$, Assumptions 1 and 2 are satisfied for $\theta=1$, $X=H^{2}_{0}(\Omega)$.
Under 2, we have the following form of a Galerkin orthogonality: Let the solution $\psi$ of (13) for given data $f\in H^{-1-\theta}(\Omega)$ satisfy $\psi\in X$. Then
$a_{h}(\psi-\psi_{h},\varphi_{h})=0\qquad\forall\varphi_{h}\in V_{h},$ (44)
where $\psi_{h}\in V_{h}$ is the solution to the discrete weak formulation of (13) satisfying
$a_{h}(\psi_{h},\varphi_{h})=\left\langle f,\varphi_{h}\right\rangle_{\Omega}\qquad\forall\varphi_{h}\in V_{h}.$
Given such bilinear form, we define the (biharmonic) Ritz projection $R_{h}\colon{X}\to V_{h},w\mapsto R_{h}w$, with $X$ from Assumption 2, given by
$a_{h}(w-R_{h}w,\chi)=0\quad\forall\chi\in V_{h}.$ (45)
3.2 Full space-time discretization
To discretize in time. we use the discontinuous Galerkin method. For that, we partition $I=(0,T]$ into subintervals $I_{m}=(t_{m-1},t_{m}]$ of length $k_{m}=t_{m}-t_{m-1}$, where $0=t_{0}<t_{1}<\dots<t_{M-1}<t_{M}=T$. The maximal and minimal time steps are denoted by $k=\max_{m}k_{m}$ and $k_{\min}=\min_{m}k_{m}$, respectively. We use the following standard notation for a function $u_{k}\in X_{k}^{r}(V)$, for a Banach space $V$, where
$X_{k}^{r}(V):=\{u_{k}\in L^{2}(I;V)\ \colon\ u_{k}\mid_{I_{m}}\in\mathbb{P}_{r}(I_{m};V),\ m=1,2,\dots,M\},$
$u_{k,m}^{+}=\lim_{\varepsilon\rightarrow 0^{+}}u_{k}(t_{m}+\varepsilon),\quad u_{k,m}^{-}=\lim_{\varepsilon\rightarrow 0^{+}}u_{k}(t_{m}-\varepsilon),\quad[u_{k}]_{m}=u_{k,m}^{+}-u_{k,m}^{-}.$ (46)
In several places of the argument we will use the following identity for any $w(t)$ with well defined one-sided limits at $t=t_{m-1}$, there holds
$([w]_{m-1},w_{m-1}^{+})_{\Omega}=\frac{1}{2}\|w_{m-1}^{+}\|^{2}_{\Omega}+\frac{1}{2}\|[w]_{m-1}\|^{2}_{\Omega}-\frac{1}{2}\|w_{m-1}^{-}\|^{2}_{\Omega}.$ (47)
We define a time projection $\pi_{k}\colon C(\bar{I})\to X_{k}^{r}(\mathbb{R})$ which on each time interval $I_{m}$ is defined by
$(\pi_{k}v-v,\phi)_{I_{m}}$ $=0\quad\forall\phi\in\mathbb{P}_{r-1}(I_{m}),\quad r>0,$ (48)
$\pi_{k}v(t_{m}^{-})$ $=v(t_{m}^{-}).$
In the case of $dG(0)$ method, the projection is defined by the second condition only. With these considerations, we can now define the fully discrete approximation to the problem (22), by introducing the fully discrete bilinear form $B_{h}\colon X_{k}^{r}(V_{h})\times X_{k}^{r}(V_{h})\to\mathbb{R}$
${B}_{h}(u_{kh},v_{kh})=$ $\sum_{m=1}^{M}(\partial_{t}\nabla u_{kh},\nabla v_{kh})_{I_{m}\times\Omega}+\int_{0}^{T}a_{h}(u_{kh},v_{kh})dt$ (49)
$+\sum_{m=2}^{M}([\nabla u_{kh}]_{m-1},{\rm NewA}v_{kh,m-1}^{+})_{\Omega}+({\rm NewA}u_{kh,0}^{+},{\rm NewA}v_{kh,0}^{+})_{\Omega}.$
Then the fully discrete problem reads: Find $\psi_{kh}\in X_{k}^{r}(V_{h})$, such that for given $f\in L^{2}(I;H^{-1-\theta}(\Omega))$, $\psi_{0}\in H^{2-\theta}_{0}(\Omega)$, it holds
${B}_{h}(\psi_{kh},\varphi_{kh})=\left\langle f,\varphi_{kh}\right\rangle_{I\times\Omega}+\left({\rm NewA}\psi_{0},{\rm NewA}\varphi_{kh,0}^{+}\right)_{\Omega},\quad\text{for all}\ \varphi_{kh}\in X_{k}^{r}(V_{h}),$ (50)
where $\theta\in[0,1]$ satisfies 1. Note that due to 1, the right hand side $\left\langle f,\varphi_{kh}\right\rangle_{I\times\Omega}$ is well defined. Moreover, we have the following parabolic Galerkin orthogonality between the continuous and fully discrete solution.
Lemma 18.
Let $\theta$ satisfy the Assumptions 1 and 2 and let $f\in L^{2}(I;H^{-1-\theta}(\Omega))$, $\psi_{0}\in H^{2-\theta}_{0}(\Omega)$. Further let $\psi$, $\psi_{kh}$ be the continuous and fully discrete solutions to (22) and (50) respectively. Then for any $\varphi_{kh}\in X^{r}_{k}(V_{h})$ there holds
$B_{h}(\psi-\psi_{kh},\varphi_{kh})=0.$ (51)
Proof.
Due to the regularity of $f$ and $\psi_{0}$, as a consequence of the interpolated regularity result in Corollary 15, we have
$\partial_{t}\Delta\psi,\mathcal{A}\psi\in L^{2}(I;H^{-1-\theta}(\Omega)).$ (52)
By 2, the regularity of $\mathcal{A}\psi$ implies $\psi\in L^{2}(I;X)$, with $X$ from Assumption 2, and thus in the definition of $B_{h}(\cdot,\cdot)$, the term $\int_{I}a_{h}(\psi,\varphi_{kh})$ is well defined. Further, as due to Theorem 11 it holds $\nabla\psi\in C(\bar{I};L^{2}(\Omega)^{2})$, all jump terms vanish and $\psi_{0}^{+}=\lim_{t\to 0+}\psi(t)=\psi(0)$. We thus obtain
${B}_{h}(\psi,\varphi_{kh})=\int_{0}^{T}\left(\partial_{t}\nabla\psi,\nabla\varphi_{kh}\right)_{\Omega}+a_{h}(\psi,\varphi_{kh})dt+\left({\rm NewA}\psi(0),{\rm NewA}\varphi_{kh,0}^{+}\right)_{\Omega}.$
By 2, due to $\psi\in L^{2}(I;X)$ we can replace the discrete elliptic bilinear form $a_{h}(\psi,\varphi_{kh})$ by the continuous elliptic operator $\left\langle\mathcal{A}\psi,\varphi_{kh}\right\rangle_{\Omega}$. Together with the definition of the Laplace operator, this yields
${B}_{h}(\psi,\varphi_{kh})=-\left\langle\partial_{t}\Delta\psi,\varphi_{kh}\right\rangle_{I\times\Omega}+\left\langle\mathcal{A}\psi,\varphi_{kh}\right\rangle_{I\times\Omega}+\left({\rm NewA}\psi(0),{\rm NewA}\varphi_{kh,0}^{+}\right)_{\Omega}.$
As by 1, $V_{h}\subset H^{1+\theta}_{0}(\Omega)$, and $H^{2}_{0}(\Omega)$ is dense in $H^{1+\theta}_{0}(\Omega)$, we can approximate $\varphi_{kh}$ by a sequence $\{\varphi^{kh}_{n}\}\subset L^{2}(I;H^{2}_{0}(\Omega))$ with $\varphi^{kh}_{n}\to\varphi_{kh}$ in $L^{2}(I;H^{1+\theta}_{0}(\Omega))$. As due to (22), it holds
$\left\langle\partial_{t}\Delta\psi,\varphi^{kh}_{n}\right\rangle_{I\times\Omega}+\left\langle\mathcal{A}\psi,\varphi^{kh}_{n}\right\rangle_{I\times\Omega}=\left\langle f,\varphi^{kh}_{n}\right\rangle_{I\times\Omega},$
the regularities of $f,\partial_{t}\Delta\psi,\mathcal{A}\psi$ imply (52) and satisfy
$\left\langle\partial_{t}\Delta\psi,\varphi_{kh}\right\rangle_{I\times\Omega}+\left\langle\mathcal{A}\psi,\varphi_{kh}\right\rangle_{I\times\Omega}=\left\langle f,\varphi_{kh}\right\rangle_{I\times\Omega}.$
Hence together with $\psi(0)=\psi_{0}$
${B}_{h}(\psi,\varphi_{kh})=\left\langle f,\varphi_{kh}\right\rangle_{I\times\Omega}+\left({\rm NewA}\psi_{0},{\rm NewA}\varphi_{kh,0}^{+}\right)_{\Omega}={B}_{h}(\psi_{kh},\varphi_{kh}),$
where the last step is precisely the definition of $\psi_{kh}$. This concludes the proof. ∎
For the fully discrete bilinear form, the following dual expression can be shown via integration by parts on each $I_{m}$ and rearranging of terms:
${B}_{h}(u_{kh},v_{kh})=$ $-\sum_{m=1}^{M}(\nabla u_{kh},\partial_{t}\nabla v_{kh})_{I_{m}\times\Omega}+\int_{0}^{T}a_{h}(u_{kh},v_{kh})\ dt$ (53)
$-\sum_{m=1}^{M-1}\left(\nabla u^{-}_{{kh},m},[{\rm NewA}v_{kh}]_{m}\right)_{\Omega}+({\rm NewA}u_{{kh},M}^{-},{\rm NewA}v_{{kh},M}^{-})_{\Omega}.$
4 Fully discrete error analysis
First, we provide the identity analogous to (47) for the symmetric form $a_{h}(\cdot,\cdot)$, namely
$a_{h}([w]_{m-1},w_{m-1}^{+})=\frac{1}{2}a_{h}(w^{+}_{m-1},w_{m-1}^{+})+\frac{1}{2}a_{h}([w]_{m-1},[w]_{m-1})-\frac{1}{2}a_{h}(w_{m-1}^{-},w_{m-1}^{-}),$ (54)
for any $w(t)\in V_{h}$. The identity follows by adding and subtracting $\frac{1}{2}a_{h}([w]_{m-1},[w]_{m-1})$ followed by direct calculations. We moreover introduce the notation
[…]
where $\Pi_{h}$ is the $H^{1}_{0}$-projection from (40).
Before we show error estimates for the discrete problem, we shall first discuss stability estimates. Analogously to the continuous results Theorems 11 and 13 we are interested in two different stability estimates. One for low regularity of data (as low as possible such that the discrete problem still is well defined), and one for higher regularity. In order to derive error estimates by a duality approach, we only require a stability estimate for higher regularity of data, which we will apply for an appropriate dual problem. This estimate can be shown under the assumptions made so far, and thus we present it first in Theorem 19, which is the discrete analog of Theorem 13. Afterwards, in Theorem 20 we shall discuss how Theorem 11 has to be generalized to the discrete case. The following proof follows the lines of [28] for second order parabolic problems.
Theorem 19.
Let $g\in L^{2}(I;H^{1}(\Omega)),\psi_{0}\in H^{1}_{0}(\Omega)$, and let $\psi_{kh}\in X^{r}_{k}(V_{h})$ be the fully discrete solution to
${B}_{h}(\psi_{kh},\varphi_{kh})=\left(\nabla g,\nabla\varphi_{kh}\right)_{I\times\Omega}+\left({\rm NewA}\psi_{0},{\rm NewA}\varphi_{kh,0}^{+}\right)_{\Omega}\quad\text{for all $\varphi_{kh}\in X_{k}^{r}(V_{h})$}.$ (56)
Then there exists a constant $C$ independent of $k$ and $h$ such that
[…]
Proof.
First we test the equation for each m=1,,Mm=1,...,M with $\varphi_{kh}|_{I_{m}}=A_{h}\psi_{kh}|_{I_{m}}$ and set $\varphi_{kh}$ to zero on all other subintervals. Then on each time subinterval $I_{m}$, we obtain
$\int_{I_{m}}({\rm NewA}\partial_{t}\psi_{kh},{\rm NewA}A_{h}\psi_{kh})_{\Omega}dt+\int_{I_{m}}\|{\rm NewA}A_{h}\psi_{kh}\|^{2}dt+({\rm NewA}[\psi_{kh}]_{m-1},{\rm NewA}A_{h}\psi_{kh,m-1}^{+})_{\Omega}=\int_{I_{m}}({\rm NewA}g,{\rm NewA}A_{h}\psi_{kh})_{\Omega}dt,$
where for $m=1$, we have made use of our convention (55). Using that
$(\partial_{t}{\rm NewA}\psi_{kh},{\rm NewA}A_{h}\psi_{kh})_{I_{m}\times\Omega}=\int_{I_{m}}\frac{1}{2}\frac{d}{dt}a_{h}(\psi_{kh},\psi_{kh})\ dt=\frac{1}{2}a_{h}(\psi_{kh,m}^{-},\psi_{kh,m}^{-})-\frac{1}{2}a_{h}(\psi_{kh,m-1}^{+},\psi_{kh,m-1}^{+})$
and since
[…]
applying the identity (54), we notice that the term with $a_{h}(\psi_{kh,m-1}^{+},\psi_{kh,m-1}^{+})$ cancels out and as a result we obtain
$\frac{1}{2}a_{h}(\psi_{kh,m}^{-},\psi_{kh,m}^{-})-\frac{1}{2}a_{h}(\psi_{kh,m-1}^{-},\psi_{kh,m-1}^{-})+\frac{1}{2}a_{h}([\psi_{kh}]_{m-1},[\psi_{kh}]_{m-1})$ $+\int_{I_{m}}\|{\rm NewA}A_{h}\psi_{kh}\|_{\Omega}^{2}dt$
$=({\rm NewA}g,{\rm NewA}A_{h}\psi_{kh})_{I_{m}\times\Omega}.$
Summing over $m$, and using the convention $\psi_{kh,0}^{-}=\Pi_{h}\psi_{0}$ yields
$\frac{1}{2}a_{h}(\psi_{kh,M}^{-},\psi_{kh,M}^{-})+\frac{1}{2}\sum_{m=1}^{M}a_{h}([\psi_{kh}]_{m-1},[\psi_{kh}]_{m-1})$ $+\int_{I}\|{\rm NewA}A_{h}\psi_{kh}\|_{\Omega}^{2}dt$
$=({\rm NewA}g,{\rm NewA}A_{h}\psi_{kh})_{I\times\Omega}+\frac{1}{2}a_{h}(\Pi_{h}\psi_{0},\Pi_{h}\psi_{0}).$
Using the Cauchy-Schwarz and Young’s inequalities, we have
[…]
Canceling, we obtain
[…]
Step 2. Time derivative terms.
Note that for the case $r=0$ there is nothing to prove here, hence we assume in this step $r\geq 1$. Noticing that $(t-t_{m-1})\partial_{t}\psi_{kh}\in\mathbb{P}_{r}(I_{m};V_{h})$, we choose $\varphi_{kh}=(t-t_{m-1})\partial_{t}\psi_{kh}$ on the time interval $I_{m}$ and zero on others. Noticing that the term with jumps vanishes, we have
$\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}\partial_{t}\psi_{kh}\|^{2}_{\Omega}dt+\int_{I_{m}}(t-t_{m-1})a_{h}(\psi_{kh},\partial_{t}\psi_{kh})dt=\int_{I_{m}}(t-t_{m-1})({\rm NewA}g,{\rm NewA}\partial_{t}\psi_{kh})_{\Omega}dt.$
For the second term, using the Cauchy-Schwarz and Young’s inequalities,
$\int_{I_{m}}(t-t_{m-1})a_{h}(\psi_{kh},\partial_{t}\psi_{kh})dt$ $=\int_{I_{m}}(t-t_{m-1})({\rm NewA}\partial_{t}\psi_{kh},{\rm NewA}A_{h}\psi_{kh})_{\Omega}dt$
$\leq\left(\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}\partial_{t}\psi_{kh}\|_{\Omega}^{2}dt\right)^{1/2}\left(\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}A_{h}\psi_{kh}\|_{\Omega}^{2}dt\right)^{1/2}$
$\leq\frac{1}{4}\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}\partial_{t}\psi_{kh}\|_{\Omega}^{2}dt+\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}A_{h}\psi_{kh}\|_{\Omega}^{2}dt.$
Similarly, Using the Cauchy-Schwarz and Young’s inequalities,
$\int_{I_{m}}(t-t_{m-1})({\rm NewA}g,{\rm NewA}\partial_{t}\psi_{kh})_{\Omega}dt$ $\leq\left(\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}g\|^{2}_{\Omega}dt\right)^{1/2}\left(\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}\partial_{t}\psi_{kh}\|^{2}_{\Omega}dt\right)^{1/2}$
$\leq\frac{1}{4}\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}\partial_{t}\psi_{kh}\|_{\Omega}^{2}dt+\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}g\|^{2}_{\Omega}dt.$
Combining estimates, and kicking back, we obtain
$\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}\partial_{t}\psi_{kh}\|^{2}_{\Omega}dt\leq C\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}g\|^{2}_{\Omega}dt+C\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}A_{h}\psi_{kh}\|^{2}_{\Omega}dt.$
Using the inverse inequality
$\int_{I_{m}}\|\chi_{k}\|^{2}_{\Omega}dt\leq Ck_{m}^{-1}\int_{I_{m}}(t-t_{m-1})\|\chi_{k}\|^{2}_{\Omega}dt$
which holds for any piecewise polynomial in time with fixed maximal degree, for example for $\chi_{k}={\rm NewA}\partial_{t}\psi_{kh}$, we have
$\int_{I_{m}}\|{\rm NewA}\partial_{t}\psi_{kh}\|^{2}_{\Omega}dt$ $\leq Ck_{m}^{-1}\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}g\|^{2}_{\Omega}dt+Ck_{m}^{-1}\int_{I_{m}}(t-t_{m-1})\|{\rm NewA}A_{h}\psi_{kh}\|_{\Omega}^{2}dt$
$\leq C\int_{I_{m}}\|{\rm NewA}g\|^{2}_{\Omega}dt+C\int_{I_{m}}\|{\rm NewA}A_{h}\psi_{kh}\|_{\Omega}^{2}dt.$
Summing over $m$ and using (57) yields
[…]
Step 3. Jump terms.
For every m=1,,Mm=1,...,M, we choose $\varphi_{kh}|_{I_{m}}=[\psi_{kh}]_{m-1}$ and set $\varphi_{kh}$ to zero on all other subintervals. Inserting this choice yields
$\int_{I_{m}}({\rm NewA}\partial_{t}\psi_{kh},[{\rm NewA}\psi_{kh}]_{m-1})_{\Omega}dt+\int_{I_{m}}({\rm NewA}A_{h}\psi_{kh},[{\rm NewA}\psi_{kh}]_{m-1})_{\Omega}dt+\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega}=\int_{I_{m}}({\rm NewA}g,[{\rm NewA}\psi_{kh}]_{m-1})_{\Omega}dt.$
Using the Cauchy-Schwarz and Young’s inequalities,
$\int_{I_{m}}({\rm NewA}\partial_{t}\psi_{kh},[{\rm NewA}\psi_{kh}]_{m-1})_{\Omega}dt$ $\leq\left(k_{m}\int_{I_{m}}\|{\rm NewA}\partial_{t}\psi_{kh}\|^{2}_{\Omega}dt\right)^{1/2}\left(k^{-1}_{m}\int_{I_{m}}\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega}dt\right)^{1/2}$
$\leq k_{m}\int_{I_{m}}\|{\rm NewA}\partial_{t}\psi_{kh}\|^{2}_{\Omega}dt+\frac{1}{4}\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega},$
where we used that $\varphi_{kh}|_{I_{m}}=[\psi_{kh}]_{m-1}$ is constant on $I_{m}$. Similarly,
$\int_{I_{m}}({\rm NewA}A_{h}\psi_{kh},[{\rm NewA}\psi_{kh}]_{m-1})_{\Omega}dt$ $\leq\left(k_{m}\int_{I_{m}}\|{\rm NewA}A_{h}\psi_{kh}\|^{2}_{\Omega}dt\right)^{1/2}\left(k^{-1}_{m}\int_{I_{m}}\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega}dt\right)^{1/2}$
$\leq k_{m}\int_{I_{m}}\|{\rm NewA}A_{h}\psi_{kh}\|^{2}_{\Omega}dt+\frac{1}{4}\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega},$
and
$\int_{I_{m}}({\rm NewA}g,[{\rm NewA}\psi_{kh}]_{m-1})_{\Omega}dt$ $\leq\left(k_{m}\int_{I_{m}}\|{\rm NewA}g\|^{2}_{\Omega}dt\right)^{1/2}\left(k^{-1}_{m}\int_{I_{m}}\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega}dt\right)^{1/2}$
$\leq k_{m}\int_{I_{m}}\|{\rm NewA}g\|^{2}_{\Omega}dt+\frac{1}{4}\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega}dt.$
Combining, we obtain
$k^{-1}_{m}\|[{\rm NewA}\psi_{kh}]_{m-1}\|^{2}_{\Omega}\leq C\left(\int_{I_{m}}\|{\rm NewA}g\|^{2}_{\Omega}dt+\int_{I_{m}}\|{\rm NewA}\partial_{t}\psi_{kh}\|^{2}_{\Omega}dt+\int_{I_{m}}\|{\rm NewA}A_{h}\psi_{kh}\psi_{kh}\|^{2}_{\Omega}dt\right).$
Summing over $m$ and using (57) and (58), we obtain
[…]
which completes the proof of the lemma. ∎
After having now shown the discrete stability result for higher regularity data, let us discuss the lower regularity setting. Note that we only include this result for completeness, and do not require it for the error estimates.
Theorem 20.
Let the 1 be satisfied, and let $f\in L^{2}(I;H^{-1-\theta}(\Omega))$, $\psi_{0}\in H^{1}_{0}(\Omega)$. Let $\psi_{kh}\in X^{r}_{k}(V_{h})$ solve (50). Then we have the following. If there exists a constant $\widetilde{C}>0$ independent of $h$, such that
[…]
then there exists a constant $C>0$ independent of $k,h$ such that:
[…]
Proof.
We begin by testing (50) with $\psi_{kh}$. For the left hand side, we obtain with the fundamental theorem of calculus, and the identity (47) applied to the jump term after canceling:
$B_{h}(\psi_{kh},\psi_{kh})$ $=\sum_{m=1}^{M}\left(\nabla\partial_{t}\psi_{kh},{\rm NewA}\psi_{kh}\right)_{I_{m}\times\Omega}+\int_{I}a_{h}(\psi_{kh},\psi_{kh})\ dt$ (60)
$\quad+\sum_{m=2}^{M}\left({\rm NewA}[\psi_{kh}]_{m-1},{\rm NewA}\psi_{kh,m-1}^{+}\right)_{\Omega}+\left({\rm NewA}\psi_{kh,0}^{+},{\rm NewA}\psi_{kh,0}^{+}\right)_{\Omega}$
[…]
$\quad+\frac{1}{2}\sum_{m=2}^{M}\left[\|{\rm NewA}\psi_{kh,m-1}^{+}\|_{\Omega}^{2}+\|{\rm NewA}[\psi_{kh}]_{m-1}\|_{\Omega}^{2}-\|{\rm NewA}\psi_{kh,m-1}^{-}\|_{\Omega}^{2}\right]+\|{\rm NewA}\psi_{kh,0}^{+}\|_{\Omega}^{2}$
[…]
For the right hand side, using (59), we obtain together with Young’s inequality
$\left\langle f,\psi_{kh}\right\rangle_{I\times\Omega}+\left({\rm NewA}\psi_{0},{\rm NewA}\psi_{kh,0}^{+}\right)_{\Omega}$ $\leq\|f\|_{L^{2}(I;H^{-1-\theta}(\Omega))}\|\psi_{kh}\|_{L^{2}(I;H^{1+\theta}(\Omega))}+\|{\rm NewA}\psi_{0}\|_{\Omega}\|{\rm NewA}\psi_{kh,0}^{+}\|_{\Omega}$
[…]
We can absorb the second and fourth term into the last line of (60) and multiply by 2, to obtain
[…]
This is almost the claimed result. If we test (50) not with $\psi_{kh}$, but with a function which coincides with $\psi_{kh}$ on $I_{1}\cup...\cup I_{m}$ and is zero on the remaining intervals, we get the final result by exactly the same argument. ∎
Remark 21.
[…]
$a_{h}(\varphi_{h},\phi_{h})=\left\langle f,\phi_{h}\right\rangle_{\Omega}\qquad\text{for all }\phi_{h}\in V_{h}.$ (61)
Remark 22.
For $C^{1}$ finite elements satisfying $V_{h}\subset H^{2}_{0}(\Omega)$ the condition (59) is trivially satisfied. For the standard $C^{0}$ finite elements used in the interior penalty method of [10] it necessarily has to hold $\theta<\frac{1}{2}$. One may then employ the arguments of [8], to obtain a localized expression of the $H^{1+\theta}_{0}(\Omega)$ norm:
$\|\varphi_{h}\|_{H^{1+\theta}(\Omega)}^{2}\leq C_{\theta}\sum_{T\in\mathcal{T}_{h}}\|\varphi_{h}\|_{H^{1+\theta}(T)}^{2}.$
[…]
We are now able to formulate our main result of the paper.
Theorem 23 (Main result).
Let the Assumptions 1 and 2 be fulfilled. Let $\psi\in\mathbb{X}$ and $\psi_{kh}\in X_{k}^{r}(V_{h})$ be the exact and the fully discrete Galerkin solutions satisfying (22) and (50) respectively for given data $f\in L^{2}(I;H^{-1-\theta}(\Omega))$, $\psi_{0}\in H^{2-\theta}_{0}(\Omega)$. Then, there exists a constant $C$ independent of $k$ and $h$ such that for any $\chi_{kh}\in X^{r}_{k}(V_{h})$
$\lVert{\rm NewA}(\psi-\psi_{kh})\rVert_{I\times\Omega}\leq C\Big{(}\lVert{\rm NewA}(\psi-\chi_{kh})\rVert_{I\times\Omega}+\lVert{\rm NewA}(R_{h}\psi-\psi)\rVert_{I\times\Omega}+\lVert{\rm NewA}(\pi_{k}\psi-\psi)\rVert_{I\times\Omega}\Big{)}.$ (62)
Proof.
Let $\chi_{kh}\in X_{k}^{r}(V_{h})$ be arbitrary but fixed. We begin by showing the estimate
$\lVert{\rm NewA}(\psi_{kh}-\chi_{kh})\rVert_{I\times\Omega}\leq C\Big{(}\lVert{\rm NewA}(v-\chi_{kh})\rVert_{I\times\Omega}+\lVert{\rm NewA}(R_{h}\psi-\chi_{kh})\rVert_{I\times\Omega}+\lVert{\rm NewA}(\pi_{k}\psi-\chi_{kh})\rVert_{I\times\Omega}\Big{)}.$ (63)
To this end we consider the dual problem
$B_{h}(\varphi_{kh},z_{kh})=({\rm NewA}(\psi_{kh}-\chi_{kh}),{\rm NewA}\varphi_{kh})_{I\times\Omega}\quad\forall\varphi_{kh}\in X_{k}^{r}(V_{h}).$ (64)
Note that due to this definition, $z_{kh}$ has a zero terminal condition at $t=T$. Hence, analogously to (55), we use the abbreviating notation $z_{kh}]_{M}=0-z_{kh,M}^{-}$. Then, due to the assumption on the data, by the Galerkin orthogonality (51) and using the dual expression of the bilinear form (53),
$\|{\rm NewA}(\psi_{kh}-\chi_{kh})\|^{2}_{I\times\Omega}$ $=B_{h}(\psi_{kh}-\chi_{kh},z_{kh})=B_{h}(\psi-\chi_{kh},z_{kh})$
$=-\sum_{m=1}^{M}\int_{I_{m}}({\rm NewA}(\psi-\chi_{kh}),{\rm NewA}\partial_{t}z_{kh})_{\Omega}\ dt+\int_{I}a_{h}(\psi-\chi_{kh},z_{kh})dt$
$\qquad-\sum_{m=1}^{M}({\rm NewA}(\psi(t_{m})-\chi_{kh}(t_{m})),[{\rm NewA}z_{kh}]_{m})_{\Omega}$
$=J_{1}+J_{2}+J_{3}.$
In the following steps, we repeatedly apply the result of Theorem 19, which was stated for a forward problem, to the dual problem (64), which is defined backwards in time. The results of the lemma still hold true with the appropriate modifications. Doing so yields for the first term the estimate
$J_{1}\leq\left(\sum_{m=1}^{M}\|{\rm NewA}(\psi-\chi_{kh})\|_{I_{m}\times\Omega}^{2}\right)^{\frac{1}{2}}\left(\sum_{m=1}^{M}\|{\rm NewA}\partial_{t}z_{kh}\|_{I_{m}\times\Omega}^{2}\right)^{\frac{1}{2}}\leq C\|{\rm NewA}(\psi-\chi_{kh})\|_{I\times\Omega}\|{\rm NewA}(\psi_{kh}-\chi_{kh})\|_{I\times\Omega}.$
To estimate $J_{2}$, we use the definition of the operators $R_{h}$ and $A_{h}$, and Theorem 19 applied to $z_{kh}$ to obtain
$J_{2}$ $=\int_{I}a_{h}(R_{h}\psi-\chi_{kh},z_{kh})dt=\int_{I}({\rm NewA}(R_{h}\psi-\chi_{kh}),{\rm NewA}A_{h}z_{kh})_{\Omega}dt$
$\leq\|{\rm NewA}(R_{h}\psi-\chi_{kh})\|_{I\times\Omega}\|{\rm NewA}A_{h}z_{kh}\|_{I\times\Omega}\leq C\|{\rm NewA}(R_{h}\psi-\chi_{kh})\|_{I\times\Omega}\|{\rm NewA}(\psi_{kh}-\chi_{kh})\|_{I\times\Omega}.$
Similarly, using the definition of the operator $\pi_{k}$, we have
$J_{3}$ $=\sum_{m=1}^{M}({\rm NewA}(\pi_{k}\psi(t_{m})-\chi_{kh}(t_{m})),[{\rm NewA}z_{kh}]_{m})_{\Omega}$
$\leq\sum_{m=1}^{M}\|{\rm NewA}(\pi_{k}\psi(t_{m})-\chi_{kh}(t_{m}))\|_{\Omega}\|[{\rm NewA}z_{kh}]_{m}\|_{\Omega}$
$\leq\left(\sum_{m=1}^{M}k_{m}\|{\rm NewA}(\pi_{k}\psi-\chi_{kh})\|^{2}_{L^{\infty}(I_{m};L^{2}(\Omega))}\right)^{1/2}\left(\sum_{m=1}^{M}k^{-1}_{m}\|[{\rm NewA}z_{kh}]_{m}\|^{2}_{\Omega}\right)^{1/2}$
$\leq C\left(\sum_{m=1}^{M}\|{\rm NewA}(\pi_{k}\psi-\chi_{kh})\|^{2}_{I_{m}\times\Omega}\right)^{1/2}\left(\sum_{m=1}^{M}k^{-1}_{m}\|[{\rm NewA}z_{kh}]_{m}\|^{2}_{\Omega}\right)^{1/2}$
$\leq C\|{\rm NewA}(\pi_{k}\psi-\chi_{kh})\|_{I\times\Omega}\|{\rm NewA}(\psi_{kh}-\chi_{kh})\|_{I\times\Omega},$
where we used one dimensional inverse inequality on $I_{m}$. Combining the estimates for $J_{1}$, $J_{2}$, and $J_{3}$ and canceling $\|{\rm NewA}(\psi_{kh}-\chi_{kh})\|_{I\times\Omega}$ on both sides, we obtain the estimate (63). From this, we obtain by inserting twice $\pm\psi$ and applying triangle inequality, that
$\lVert{\rm NewA}(\psi-\psi_{{kh}})\rVert_{I\times\Omega}\leq C\Big{(}\lVert{\rm NewA}(\psi-\chi_{kh})\rVert_{I\times\Omega}+\lVert{\rm NewA}(R_{h}\psi-\psi)\rVert_{I\times\Omega}+\lVert{\rm NewA}(\pi_{k}\psi-\psi)\rVert_{I\times\Omega}+2\lVert{\rm NewA}(\psi-\chi_{kh})\rVert_{I\times\Omega}\Big{)}.$
Collecting the $\lVert{\rm NewA}(\psi-\chi_{kh})\rVert_{I\times\Omega}$ terms together then concludes the proof. ∎
4.1 Specific spacial finite element spaces
Here, we provide several examples of the finite element spaces for solving stationary biharmonic problem.
4.1.1 Conformal $C^{1}$ elements
It is natural to seek a finite element space $V_{h}\subset H^{2}_{0}(\Omega)$, i.e. a conformal finite element approximation. In two dimensions, there are a number of such constructions, usually require rather high polynomial order, we will focus only on $C^{1}$ Argyris elements, ( cf. [11, p. 273] for other finite element methods for fourth order problems). Trivially, 1 is satisfied with $\theta=1$ and 2 holds since the method is conformal and the discrete bilinear form is coincides with the continuous one, see Remark 17. In the case of $\Omega$ being convex, the standard duality argument and a priori error estimates show
$\|\psi-R_{h}\psi\|_{H^{1}(\Omega)}\leq Ch^{2}\|\psi\|_{H^{3}(\Omega)}.$ (65)
4.1.2 The $C^{0}$ interior penalty method
This method is very attractive since it uses Lagrange elements of degree $l\geq 2$. To define the method, we need some additional notation. Let $\mathcal{E}_{h}$ be the set of edges in $\mathcal{T}_{h}$. For $e\in\mathcal{E}_{h}$ and $v\in(X+V_{h})$, with $X$ from Assumption 2, we define the jump $\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]$ of the normal derivative of $v$ across an edge $e$ and the average of the second normal derivative $\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}$ of $v$ on an edge $e$ as follows. If $e\subset\Omega$, we take $n_{e}$ to be one of the two unit vectors normal to $e$. Then $e$ is the common side of two triangles $T_{+}\in\mathcal{T}_{h}$ and $T_{-}\in\mathcal{T}_{h}$, where $n_{e}$ is pointing from $T_{-}$ to $T_{+}$. Thus, on such $e$ we define
$\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]=\frac{\partial v_{T_{+}}}{\partial n}\mid_{e}-\frac{\partial v_{T-}}{\partial n}\mid_{e}\quad\text{and}\quad\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}=\frac{1}{2}\left(\frac{\partial^{2}v_{T_{+}}}{\partial n^{2}}\mid_{e}+\frac{\partial^{2}v_{T-}}{\partial n^{2}}\mid_{e}\right).$
We note that the above definitions do not depend on the choice of $n_{e}$. If $e\subset\partial\Omega$, we take $n_{e}$ to be the unit normal pointing outside $\Omega$ and
$\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]=-\frac{\partial v}{\partial n_{e}}\quad\text{and}\quad\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}=-\frac{\partial^{2}v}{\partial n_{e}^{2}}.$
We define the bilinear form $a_{h}:(X+V_{h})\times(X+V_{h})\to\mathbb{R}$ by
$a_{h}(v,w)$ $=\sum_{T\in\mathcal{T}_{h}}\int_{T}D^{2}v:D^{2}w\,dx$
$+\sum_{e\in\mathcal{E}_{h}}\int_{e}\Big{(}\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}v}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}\left[\hskip-2.52971pt\left[\frac{\partial w}{\partial n}\right]\hskip-2.52971pt\right]+\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]\left\{\hskip-3.97432pt\left\{\frac{\partial^{2}w}{\partial n^{2}}\right\}\hskip-3.97432pt\right\}+\frac{\eta}{|e|}\left[\hskip-2.52971pt\left[\frac{\partial v}{\partial n}\right]\hskip-2.52971pt\right]\left[\hskip-2.52971pt\left[\frac{\partial w}{\partial n}\right]\hskip-2.52971pt\right]\Big{)}ds,$
where
$D^{2}v:D^{2}w=\sum_{i,j=1}^{2}\partial^{2}_{ij}v\ \partial^{2}_{ij}w,$
$|e|$ is the length of the edge $e$ and $\eta>0$ is a sufficiently large penalty constant. For this choice of discretization, the results of [10, Lemma 5] show that for any polygonal domain $\Omega$, the 2 holds for $\theta=\frac{1}{2}-\varepsilon$, where $\varepsilon>0$ can be chosen arbitrarily small. It is a standard result, that the 1 is also satisfied. The space $X$ in this case is given by $H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}-\varepsilon}(\Omega)$. Hence, the Ritz projection $R_{h}$ introduced in (45) is well defined for any $\psi\in H^{2}_{0}(\Omega)\cap H^{\frac{5}{2}-\varepsilon}(\Omega)$. Moreover, if $\Omega$ is convex, we have the following error estimate (cf. [10], Theorem 5, with $\alpha=1$)
$\|\psi-R_{h}\psi\|_{H^{1}(\Omega)}\leq Ch^{2}\|\psi\|_{H^{3}(\Omega)}.$ (66)
Thus, in the above setting setting, for conformal $C^{1}$ and $C^{0}$ interior penalty methods, we obtain the following rates of convergence.
Theorem 24.
Let $\Omega$ be convex, $f\in L^{2}(I;H^{-1}(\Omega))$ and $\psi_{0}\in H^{2}_{0}(\Omega)$. Let further $\psi$, $\psi_{kh}$ solve the continuous and fully discrete time-dependent stream-function equations (22) and (50), respectevely. Then there exists $C>0$ independent of $h$ and $k$ such that
$\lVert\nabla(\psi-\psi_{kh})\rVert_{L^{2}(I;L^{2}(\Omega))}\leq C(k+h^{2})\left(\lVert f\rVert_{L^{2}(I;H^{-1}(\Omega))}+\lVert\psi_{0}\rVert_{H^{2}_{0}(\Omega)}\right).$
Proof.
Due to the discussion previous to the theorem, the Assumptions 1 and 2 are fulfilled. Hence we need to estimate the terms presented in Theorem 23. By the regularity results derived in Theorem 13, the stationary error estimate of the Ritz projection (66) and the error estimate of [27, Lemma 5.13] for the time projection $\pi_{k}$,we have
$\lVert{\rm NewA}(\pi_{k}\psi-\psi)\rVert_{L^{2}(I\times\Omega)}\leq Ck\lVert\partial_{t}{\rm NewA}\psi\rVert_{L^{2}(I\times\Omega)}\leq Ck\left(\lVert f\rVert_{L^{2}(I;H^{-1}(\Omega))}+\lVert\psi_{0}\rVert_{H^{2}_{0}(\Omega)}\right),$ (67)
$\lVert{\rm NewA}(R_{h}\psi-\psi)\rVert_{L^{2}(I\times\Omega)}\leq Ch^{2}\lVert\psi\rVert_{L^{2}(I;H^{3}(\Omega))}\leq Ch^{2}\left(\lVert f\rVert_{L^{2}(I;H^{-1}(\Omega))}+\lVert\psi_{0}\rVert_{H^{2}_{0}(\Omega)}\right).$ (68)
In the best approximation term, following the proof of [25, Theorem 12] we choose $\chi_{kh}=\Pi_{h}\pi_{k}\psi$, where $\Pi_{h}$ is the $H^{1}_{0}$ projection introduced in (40). This, together with triangle inequality and the stability of $\Pi_{h}$ in $H^{1}_{0}(\Omega)$, see (41), yields
$\lVert{\rm NewA}(\psi-\chi_{kh})\rVert_{L^{2}(I\times\Omega)}$ $\leq\lVert{\rm NewA}(\psi-\Pi_{h}\psi)\rVert_{L^{2}(I\times\Omega)}+\lVert{\rm NewA}(\Pi_{h}(\psi-\pi_{k}\psi))\rVert_{L^{2}(I\times\Omega)}$
$\leq\lVert{\rm NewA}(\psi-\Pi_{h}\psi)\rVert_{L^{2}(I\times\Omega)}+\lVert{\rm NewA}(\psi-\pi_{k}\psi)\rVert_{L^{2}(I\times\Omega)}.$
We have already estimated the second term in (67), and it is a classical result from numerical analysis of the Poisson problem, that due to $l\geq 2$ and the regularity of $\psi$, the projection $\Pi_{h}$ satisfies the error estimate
$\lVert{\rm NewA}(\psi-\Pi_{h}\psi)\rVert_{L^{2}(I\times\Omega)}\leq Ch^{2}\lVert\psi\rVert_{L^{2}(I;H^{3}(\Omega))}\leq Ch^{2}(\lVert f\rVert_{L^{2}(I;H^{-1}(\Omega))}+\lVert\psi_{0}\rVert_{H^{2}_{0}(\Omega)}),$
which concludes the proof. ∎
Naturally, $\operatorname{\mathbf{curl}}(\psi_{kh})=(\partial_{2}\psi_{kh},-\partial_{1}\psi_{kh})^{T}$ is discontinuous across triangulation boundaries. One way to address this issue is postprocessing [10], but this topic lies outside of the scope of this paper.
5 Numerical results
To demonstrate the theoretical error estimates derived in this work, we generate an example where the analytic solution $\mathbf{u}$ to the time-dependent Stokes equations is known. We choose $I=(0,1]$, $\Omega=[0,1]^{2}$ and calculate $\mathbf{w}=\operatorname{\mathbf{curl}}(\Phi)$, where $\Phi\colon\mathbb{R}^{2}\to\mathbb{R},\Phi(x)=\sin(2\pi x_{1})^{2}\sin(2\pi x_{2})^{2}$. We then choose $\mathbf{u}=\sin(2\pi t)\mathbf{w}$ and set
$\mathbf{f}=\partial_{t}\mathbf{u}-\Delta\mathbf{u}.$
By construction $\mathbf{u}$ satisfies $\nabla\cdot\mathbf{u}=0$, $\mathbf{u}(0)=\mathbf{0}$ and $\mathbf{u}|_{\partial\Omega}=0$, and thus solves the Stokes equations with right hand side $\mathbf{f}$, together with an associated pressure $p=0$. We first discretize the standard weak formulation of these equations by dG(0) in time and the MINI element in space. The numerically observed errors for this setting are reported in Figure 1. We then proceed by modifying the right hand side $\mathbf{f}$ by a gradient field, i.e., set the new right hand side to
$\tilde{\mathbf{f}}=\mathbf{f}+10^{5}\begin{pmatrix}x_{1}^{-0.49}\\ 0\end{pmatrix}$
As $\tilde{\mathbf{f}}$ and $\mathbf{f}$ only differ by a gradient field, they yield the same velocity solution $\mathbf{u}$, and only the pressure changes to $\tilde{p}=p+10^{5}(1.49\cdot x_{1}^{1.49})^{-1}$. As for the discretization of the Stokes with mixed finite elements, the velocity error is bounded by the sum of best approximation errors for velocity and pressure, the error in the velocity increases, which is reported in Figure 2. Lastly we set up a discrete stream-function formulation by taking $g=-\operatorname{curl}(\mathbf{f})$ and solving
$B_{h}(\psi_{kh},\varphi_{kh})=\left(g,\varphi_{kh}\right)_{I\times\Omega}\quad\forall\varphi_{kh}\in X^{r}_{k}(V_{h}).$
We then report the error $\mathbf{u}-\operatorname{\mathbf{curl}}(\psi_{kh})$ in Figure 3. Note that we observe exactly the error estimates derived in Theorem 24. As in the setting of this problem, $\operatorname{curl}$ is applied to the right hand side $\mathbf{f}$ or $\tilde{\mathbf{f}}$ respectively, the same convergence behaviour is obtained for the modified right hand side $\tilde{\mathbf{f}}$, showing the pressure robustness of this method.
Fig. 1: Convergence for discretization with MINI element
Fig. 2: Convergence for discretization with MINI element for Gradient field
Fig. 3: Convergence for Stream-function formulation